### v2

# Massively Multilingual Adaptation of Large Language Models Using Bilingual Translation Data

Abstract
This paper investigates a critical design decision in the practice of massively multilingual continual pre-training — the inclusion of parallel data. Specifically, we study the impact of bilingual translation data for massively multilingual language adaptation of the Llama3 family of models to 500 languages. To this end, we construct the MaLA bilingual translation corpus, containing data from more than 2,500 language pairs. Subsequently, we develop the EMMA-500 Llama 3 suite of four massively multilingual models — continually pre-trained from the Llama 3 family of base models extensively on diverse data mixes up to 671B tokens — and explore the effect of continual pre-training with or without bilingual translation data. Comprehensive evaluation across 7 tasks and 12 benchmarks demonstrates that bilingual data tends to enhance language transfer and performance, particularly for low-resource languages. We open-source the MaLA corpus, EMMA-500 Llama 3 suite artefacts, code, and model generations.
Website: mala-lm.github.io/emma-500-gen2
[Uncaptioned image] Models: huggingface.co/collections/MaLA-LM $\rightarrow$ EMMA-500
[Uncaptioned image] Data: huggingface.co/datasets/MaLA-LM/mala-bilingual-translation-corpus
[Uncaptioned image] Evaluation: github.com/MaLA-LM/emma-500
1 Introduction
Large language models (LLMs) pre-trained on massive data have promoted multilingual natural language processing (NLP). However, multilingual models such as BLOOM (78) and Llama (91, 92) often struggle with low-resource languages and are still limited in their language coverage (32, 80, 33). Recent works extend pre-trained LLMs into multiple languages via continual pre-training (CPT). For example, LLaMAX (54) and xLLMs-100 (44) adopt CPT and instruction fine-tuning to extend existing LLMs into 100 languages, and MaLA-500 (51) and EMMA-500 (34) perform continual pre-training (low-rank and full-parameter CPT using Llama 2) to adapt LLMs into 500 languages. Despite these efforts, challenges still remain in adapting LLMs to low-resource languages, especially in a massively multilingual scenario with more than 500 languages.
This paper studies CPT in a massively multilingual setting. Prior work like LlaMAX (54) uses both monolingual and parallel texts for CPT in 100 languages, and EMMA-500 Llama 2 (34) uses only monolingual texts in 500 languages. Our novelty is to scale the number of languages up to more than 500 languages, compile a new corpus with bilingual translation data (MaLA), and study the comparative effects of continual pre-training with monolingual and bilingual translation data. [^1] Monolingual data consists of texts written in a single language. Bilingual translation data, also called parallel corpora, comprises pairs of sentences in two different languages that express the same meaning. In this paper, we treat the terms bilingual translation corpora/texts/data, bilingual corpora/texts/data, parallel corpora/texts/data, and bitexts as equivalent.
Contributions
Our contributions are three-fold:
  • Data: We compile a bilingual translation corpus for Massive Language Adaptation in more than 2,500 language pairs and 500 languages, namely the MaLA translation corpus.
  • Models: We train and release 4 models, namely EMMA-500 Llama 3/3.1 Mono/Bi for Enhancing Massively Multilingual Adaptation, [^2] We use MaLA and EMMA-500 to name the corpus and models, following the naming convention of EMMA-500 Llama 2 (34) which is not an artifact of this paper. MaLA and EMMA-500 are a collection of corpora and models. In this paper, we compile and release the MaLA bilingual translation corpus, i.e., [Uncaptioned image] MaLA-LM/mala-bilingual-translation-corpus. “Mono” and “Bi” in model names indicate CPT on monolingual (fig. 2(b)) and bilingual (fig. 2(a)) mixes, respectively. by continually pre-training of Llama 3 & 3.1 (8B) using both monolingual and bilingual MaLA corpus augmented with diverse data types, up to 671B tokens.
  • Evaluation: We conduct a comprehensive evaluation across 7 tasks and 12 benchmarks. Our empirical evaluation ablates the impact of two diverse data mixes and analyzes gains in task generalization and multilingual robustness.
Evaluation Results and Findings
  • CPT using a data mix with bilingual translation data generally exhibits better multilingual performance than a monolingual mix [^3] A monolingual mix (fig. 2(b)) contains monolingual data in different languages but not in the aligned format as parallel data., particularly in low-resource languages and in machine translation tasks that directly benefit from parallel data.
  • Heavily pre-trained models (e.g., Llama 3 and 3.1) that consume more training tokens are more resistant to further adaptation than English-centric models (e.g., Llama 2) when scaling to include many additional languages.
  • As for overall performance, our EMMA-500 models are the best at machine translation (Flores200) and competitive at text classification (Taxi1500 and SIB-200) and commonsense reasoning (XCOPA and XStoryCloze).
  • EMMA-500 CPT models exhibit a lower average accuracy on the BELEBLE comprehension benchmark, but they outperform baselines across a greater number of languages.
While multilingual models can achieve broad coverage, perfect uniformity across all tasks and languages remains an unattainable goal. However, we show that multilingual performance and language equality can be pushed forward with parallel training data.
Outline
Section˜2 presents the data and model training with a newly compiled bilingual translation corpus introduced in Section˜2.1, data mixing introduced in Section˜2.2, and settings for model training introduced in Section˜2.3. Appendices˜B and C describe the details about the MaLA translation corpus and data mixes. We evaluate the resulting models and discuss the evaluation results in Section˜3. Detailed evaluation setup and per-benchmark and per-language results are presented in Appendices˜D and E respectively. We conclude the paper in Section˜5. Sections˜4 and F introduce related work and ethics consideration.
2 Data and Model Training
2.1 MaLA Translation Corpus
Bilingual translation corpora are language datasets that contain text data in one language aligned with text data in another language. We extend the MaLA (Massive Language Adaptation) corpus (34) by incorporating parallel data in more than 500 languages and 2,500 language pairs. The resulting parallel dataset is named the MaLA translation corpus (MaLA for short), which is suitable for adapting language models in massively multilingual scenarios. The section describes the process of building the MaLA translation corpus in a way similar to the MaLA corpus with monolingual texts, but focuses on bilingual texts.
We follow a similar data integration pipeline to the process of compiling the MaLA corpus (34), including extraction, harmonization, language code normalization, and writing system recognition. The bilingual data comes from various sources, including OPUS (89), NLLB (64), and Tatoeba (90). The datasets from OPUS are made by an existing compilation: Lego-MT (104). Table˜8 in Section˜B.1 shows the data sources for bilingual texts. The main difference in the script recognition, as well as language code conversion, with the bilingual corpora, is in the form of the label; we obtain a label in the form of a language pair, e.g., eng_Latn-zho_Hani.
Language Code Normalization
Language code normalization converts various language codes into a standardized format to ensure consistency and compatibility across different systems and applications. With the bilingual corpora, we face similar issues as with the monolingual ones when converting language denotations given in OPUS [^4] https://github.com/Helsinki-NLP/OPUS to ISO 639-3 language codes. Moreover, with bilingual corpora, we want to specify dialects based on the ISO 3166-1 alpha-3 standard. The recognition and handling of language codes are based on the following procedure:
  • If the language code of a dataset provided by OPUS matches a language code in ISO-639-3, then we consider it such.
  • If the language code does not match one in ISO-639-3, we use the langcodes package [^5] https://pypi.org/project/langcodes/ to convert it to ISO-639-3.
  • If the above steps fail, we assign “unknown” as the language code.
Writing System Recognition
We implement writing system identification following ISO 15924 standards using the GlotScript library (38). For script detection, we analyze 100-line samples by default, reverting to first-line analysis when standard detection fails—a known limitation affecting both our bilingual and monolingual corpora (34). We do not classify a dataset into multiple scripts, even in cases of code-mixing, where multiple scripts are used.
Data Cleaning
The bilingual corpus compilation faces significant quality variability, such as noisy sentence pairs. The Lego-MT dataset (104) applies some data cleaning, including deduplication, removing missing translations, and length mismatching. For other data sources, we add various procedures, including dataset-specific cleaning and deduplication, after integrating all data sets into our collection. We eliminate lines that contain the exact word or character repeated more than five consecutive times. This problem appears in particular in the Tatoeba parallel training data (90) and in the majority of these cases is erroneous. We use OpusFilter (8) for deduplication of data points.
Key Statistics
After pre-processing and cleaning, we obtain the MaLA translation corpus in 2,507 language pairs. Table˜1 shows the total number of whitespace-separated tokens and the number of language pairs across different resource categories. Compared with Lego-MT and NLLB, MaLA has a similar number of language pairs but more tokens. We categorize language pairs into 5 resource levels based on token counts: high-resource (>1B), medium-high (>500M), medium (>100M), medium-low (>10M), and low (>1M). Different from the monolingual MaLA, we add two categories of “very high” and “very low” resources in the resource level classification, i.e., very high-resource pairs (>10B) and very low-resource (<1M). In total, there are more than 426B tokens in the MaLA translation corpus. [^6] Note that whitespace-based token counting is not accurate for a language where words are not typically separated by spaces. In these languages, the absence of whitespace makes it challenging to determine token boundaries, leading to inaccurate token counts when using whitespace as a delimiter. We use whitespace as the delimiter because of its efficiency in processing text. We further sample the MaLA translation corpus for continued training LLMs, considering a balanced corpus size and language coverage.
Table 1: Key statistics of the MaLA translation corpus and comparison with existing parallel corpora.
Categories Lego-MT NLLB MaLA
Pairs Tokens Pairs Tokens Pairs Tokens
very high 4 5.1E+10 4 2.8E+10 4 8.5E+10
high 51 1.4E+11 51 1.2E+11 83 2.1E+11
medium-high 22 1.5E+10 65 4.5E+10 67 4.7E+10
medium 75 1.8E+10 264 6.1E+10 281 6.4E+10
medium-low 113 4.1E+09 480 2.0E+10 508 2.0E+10
low 350 1.4E+09 491 1.7E+09 655 2.5E+09
very low 1893 1.7E+08 1154 1.3E+08 909 1.8E+08
sum 2508 2.2E+11 2509 2.8E+11 2507 4.3E+11
2.2 Data Mixing
We blend the data compiled in MaLA with multilingual non-parallel data obtained from the cleaned and deduplicated MaLA corpus (34) along with texts selectively sourced from factual and high-quality domains with a view to retaining the knowledge acquired during the annealing phase of pre-training (31). We mirror existing practice and sample books and scientific papers (82) along with instruction-like (58) data. Following Ji et al. (34), we use scientific papers and books from CSL (47), pes2o (81), and free e-books from the Gutenberg project [^7] https://www.gutenberg.org/ compiled by Faysse (22). Multilingual instruction data is sourced from the training set of xp3x (Crosslingual Public Pool of Prompts eXtended) [^8] https://hf.co/datasets/CohereForAI/xP3x and the Aya collection [^9] https://hf.co/datasets/CohereForAI/aya_collection_language_split. Finally, we augment our mixes with code and code-adjacent procedural text due to its demonstrable benefits towards reasoning and entity-tracking (77, 69) (details in Section˜C.1).
We manually mix up different types of data to balance the language coverage across different levels of resources and types while ensuring that low- and medium-resource languages remain overrepresented. Figure˜2 shows the composition categorized by language resources according to the number of tokens. Notably, medium-resource language pairs contribute the majority of bilingual data, and medium-high-resource languages are the largest category for monolingual and instruction data.
(a) Bilingual translation data
(b) Monolingual data
(c) Instruction data
Figure 1: Composition by language resource levels of different data types.
Two Data Mixes
We make two data mixes to ablate the effect of incorporating bilingual texts into continual pre-training. The first data mix is a bilingual mix (Figure˜2(b)), which incorporates various data types. The second is a monolingual mix (Figure˜2(b)), which is derived from the bilingual mix but specifically omits any bilingual data, focusing solely on a subset with monolingual texts per document. Detailed data statistics of these two mixes are presented in Table˜10.
Most translation data are aligned sentences, and there is not sufficient document-level aligned data publicly available to be collected in the MaLA corpus, especially in the massively multilingual scenario. When using translation data, we concatenate the source and target language texts in a specific format to form a chunk of pairs in the same language pair. For every ten samples, i.e., sentence pairs, to make a document for training, [^10] The choice of 10 is inspired by “yí mù shí háng”, a Chinese idiom that literally translates to “one glance ten lines”, means that someone reads very quickly and efficiently. the format is structured as follows:
[…]
# 8 lines for 8 samples
[{src_lang_code}]: {src_text} [{tgt_lang_code}]: {tgt_text}
In this format, the notation $\{\cdot\}$ denotes the variables for source and target language codes and texts. This method allows us to work with pseudo-document-level data and clearly delineate between the source and target languages, facilitating better processing and understanding of the translation data without switching between multiple languages within the context window. By organizing the data this way, we facilitate the model to learn from the relationships between the two languages effectively. This structured approach ensures clarity and consistency in how bilingual data is presented, making it easier to process and analyze.
Figure˜2 shows the composition of our two data mixes. In the monolingual mix, monolingual web-crawled text is the largest data type, as its name suggests. The bilingual mix incorporates a considerably bigger portion of bilingual texts, 6% more than the monolingual texts. Continual training on these two mixes facilitates the adaptation of LLMs to massively multilingual languages and analyzes the effect of scaling massively multilingual training using bilingual data.
(a) Data mix 1: bilingual
(b) Data mix 2: monolingual
Figure 2: Two date mixes and their composition. The bilingual mix includes all types of data. The monolingual mix consists of a subset of the bilingual mix that excludes bilingual data.
Manual Curation of Data Mixes
Our data mixes in Figure˜2 are crafted through experience and intuition. We draw on our knowledge and understanding of the domain to create effective data combinations to balance the language resources and represent different text types. This allows for a reasonable selection that can capture the complexities of the data landscape and is aligned with the goals for massively multilingual adaptation. From an algorithmic perspective, grid search or other similar methods could be used to explore the space of possible data distributions. However, algorithmic search involves systematically exploring a predefined set of hyperparameter values by evaluating all possible combinations through model training on the searched mixes, which can become computationally expensive and time-consuming, especially as the number of parameters increases. This makes it challenging to apply algorithms like grid search effectively in scenarios where the data mix needs to be optimized, which would require thousands of training runs and tens of millions of GPU hours or even more. Searching for an optimal (or near-optimal) data mix is practically infeasible. Our paper focuses on offering a useful resource for continual pre-training in a massively multilingual scenario. To do this, we select the corpus to maintain diversity and balance among various languages and text types that align with the model’s intended goal.
2.3 Model Training
We continue training the decoder-only Llama 3 models (8B parameters) using the causal language modeling objective, exposing the pre-trained model to new data and languages to develop our EMMA-500 model. To enhance efficiency, we use training strategies that optimize memory usage, precision handling, and distributed training. EMMA-500 Llama 3 series models are trained on the LUMI Supercomputer, powered by 100% renewable and carbon-neutral energy, utilizing 64 compute nodes with 256 AMD MI250x GPUs (512 AMD Graphics Compute Dies) with the GPT-NeoX framework (6).
We continue training the base models in a full-parameter manner without modifying the tokenizer. The training setup includes a global batch size of 2048 and sequence lengths of 8192 tokens. Model- and data-specific settings are presented in Table˜2. For training on the monolingual mix, the process spans 25,000 steps, accumulating a total of 419 billion Llama 3 tokens. For the bilingual mix that contains more tokens, we train for more steps, up to 40,000, leading to a total of 671 billion tokens. We employ the Adam optimizer (39) with a learning rate of 0.0001, betas set to [0.9, 0.95], and an epsilon of 1e-8. We experiment with different learning rates and evaluate early checkpoints trained up to 5,000 steps (84 billion tokens). We find that training with the original learning rate of 0.0003 used by the Llama 3 leads to instability, such as many fast spikes, resulting in poor performance. Thus, we opt for a smaller learning rate with a more stable training curve. All experiments consume more than 800k GPU hours on the LUMI Supercomputer. However, we could not perform a grid search on the learning rate due to the constraint of computing resources.
A cosine learning rate scheduler, with a warm-up of 1,000 and 2,000 iterations for monolingual and bilingual mixes, respectively, is used to regulate learning dynamics. To optimize memory consumption, activation checkpointing is applied. Additionally, we leverage mixed-precision techniques, utilizing bfloat16 for computations while maintaining FP32 for gradient accumulation to balance efficiency and accuracy.
Table 2: Continual pre-trained models and settings.
Base Model Data Mix Our Model Name Steps Warmup Tokens
Llama 3 Monolingual (fig.˜2(b)) EMMA-500 Llama 3 Mono 25,000 1,000 419B
Bilingual (fig.˜2(a)) EMMA-500 Llama 3 Bi 40,000 2,000 671B
Llama 3.1 Monolingual (fig.˜2(b)) EMMA-500 Llama 3.1 Mono 25,000 1,000 419B
Bilingual (fig.˜2(a)) EMMA-500 Llama 3.1 Bi 40,000 2,000 671B
3 Evaluation and Discussion
This section is structured to systematically evaluate the overall performance of our models on multilingual and bilingual benchmarks, assessing improvements in both text understanding and generation. We analyze the impact of bilingual continual pre-training on multilingual language models and conduct ablation studies to isolate the contributions of bilingual pre-training compared to monolingual training. We also focus on low-resource language performance, demonstrating how continual pre-training enhances representation and generalization for underrepresented languages. No single model can be universally the best among all baselines across the full spectrum of multilingual tasks, benchmarks, and languages. We provide an analysis of language gains and failure cases, identifying which languages benefit the most and highlighting remaining challenges.
Tasks and Benchmarks
We evaluate all models on 7 tasks and 12 benchmarks that cover from 10 to 1500 languages. Table˜11 shows the details of those tasks, benchmarks, evaluation metrics, the number of languages and samples per language, and the domain. We do not use LLMs-as-a-judge (46) for evaluation due to its well-known limitations, especially in multilingual scenarios, such as low agreement with human judges (18).
Baselines
We consider open-weight decoder-only models with 7-9B parameters as baselines. We primarily compare our CPT models with the original Llama 3/3.1 base models (20) and the LlaMAX models (54) continually trained from Llama 3. We also compare with a wide set of baselines (Section˜D.3), including (1) Llama 2 models and their CPT models; (2) multilingual models, including recent advances such as Aya 23 (7), Gemma (88), Qwen (103), and Marco-LLM (61).
3.1 Impact of Mono- vs. Bilingual Training
To understand the role of continual training with bilingual translation data, we conduct controlled ablation studies by comparing monolingual continual pre-training (with the monolingual corpus only, i.e., data mix 2 in Figure˜2(b)) vs. bilingual continual pre-training (with monolingual and bilingual extension, i.e., data mix 1 in Figure˜2(a)). Figure˜3 compares their average performance on each benchmark.
For CPT with Llama 3 as shown in Figure˜3(a), continual training on data mix with bilingual translation data consistently improves commonsense reasoning, natural language inference, reading comprehension, machine translation, and sometimes improves text classification when evaluated on Taxi1500, and retains similar summarization and math performance to the original base models.
For CPT with Llama 3.1 as shown in Figure˜3(b), continual training on data mix with bilingual translation data consistently improves text classification, commonsense reasoning, and machine translation. Except for summarization tasks, our CPT models trained on the monolingual mix are usually better than on the bilingual mix. More remarkably, CPT with both Llama 3 and 3.1 shows a large improvement on machine translation with an increase from 9% to 140% in terms of BLEU or chrF++ scores on translation directions from and to English on the Flores200 dataset.
Our study provided insights into how bilingual texts contribute to language adaptation, transferability, and performance stability. Overall, continual training with bilingual translation data tends to improve the multilingual performance, especially for machine translation, which benefits directly from the use of parallel texts in training.
(a) Llama 3
(b) Llama 3.1
Figure 3: Comparison of monolingual and bilingual CPT. The scores are averaged across all evaluated languages of the corresponding benchmarks. The baseline model LlaMAX does not have a CPT variant trained on Llama 3.1. Our models show a tendency for bilingual CPT to be better than monolingual CPT in most benchmarks and a remarkable advance on the Flores200 translation benchmark.
3.2 Low-Resource Language Performance
Low-resource languages often struggle with data scarcity and representation biases in multilingual models. We evaluate how baselines and our models perform on low-resource languages, and analyze whether continual pre-training on our data mixtures, especially the bilingual extension, can mitigate these limitations.
We focus on two benchmarks, SIB-200 and Flores200, with more than 200 languages, and one benchmark, Taxi1500, with more than 1500 languages, and examine the low-resource languages of these benchmarks according to the categorization in Table˜9. Table˜3 shows the average performance on low-resource languages. Our EMMA-500 Llama 3 bilingual model obtained the best translation performance, followed by our Llama 3.1 bilingual one. They surpass the advanced Macro-LLM by a large margin on English to other translations and a small margin on other to English translations. For text classification, our models experience different levels of drops in the low-resource languages of SIB-200. Notably, our bilingual models perform the best on Taxi1500, showing their outstanding performance on low-resource languages validated on this massively multilingual benchmark. The results on low-resource languages show that CPT on massive parallel data enhances the performance on low-resource languages, especially for massively multilingual classification and translation.
Table 3: Performance on low-resource languages. The text classification task uses SIB-200 (SIB) and Taxi1500 (Taxi) datasets. Underline and bold represents the absolute best, underline means the second best, and bold signifies the best within a specific group. Our EMMA-500 models trained on bilingual data are the best two models in most cases.
Model SIB Taxi Flores200 (Eng-X) Flores200 (X-Eng)
ACC ACC chrF++ BLEU chrF++ BLEU
Llama 2 7B 22.74 18.04 16.23 4.93 31.26 13.69
Llama 2 7B Chat 26.02 16.01 18.12 5.30 32.61 13.05
CodeLlama 2 7B 23.84 17.15 15.98 4.57 29.47 11.52
LLaMAX Llama 2 7B 10.78 23.53 7.92 0.86 14.08 2.08
LLaMAX Llama 2 7B Alpaca 28.66 15.57 30.45 13.36 44.08 23.85
MaLA-500 Llama 2 10B v1 23.13 24.76 6.64 0.65 13.89 2.55
MaLA-500 Llama 2 10B v2 19.04 22.72 6.93 0.55 15.76 3.17
YaYi Llama 2 7B 25.03 17.98 16.04 4.81 32.30 13.76
TowerBase Llama 2 7B 19.53 18.06 17.15 5.10 32.44 14.50
TowerInstruct Llama 2 7B 20.81 17.92 16.84 3.54 26.32 5.24
EMMA-500 Llama 2 7B 31.97 21.73 35.88 16.87 48.00 27.34
Occiglot Mistral 7B v0.1 33.27 22.62 17.23 4.69 32.13 14.00
Occiglot Mistral 7B v0.1 Instruct 34.87 19.59 16.92 4.31 32.61 12.31
BLOOM 7B 17.73 14.93 12.32 2.67 28.51 9.96
BLOOMZ 7B 29.99 17.00 16.83 7.48 35.63 21.06
YaYi 7B 36.40 16.19 14.20 4.23 21.25 4.72
Aya 23 8B 42.20 22.52 16.89 6.41 33.31 14.48
Aya Expanse 8B 58.06 19.08 25.13 6.78 37.85 13.61
Gemma 7B 59.97 16.55 24.75 9.56 45.65 25.47
Gemma 2 9B 47.19 21.48 28.50 12.71 41.36 25.19
Qwen 1.5 7B 48.94 8.18 18.87 5.93 37.01 16.42
Qwen 2 7B 56.06 23.01 18.17 5.57 38.90 18.40
Qwen 2.5 7B 54.85 17.82 18.46 5.71 40.11 19.97
Marco-LLM GLO 7B 65.07 21.97 24.79 9.51 45.89 26.38
Llama 3 8B 65.53 23.61 25.84 10.51 45.72 25.49
Llama 3.1 8B 63.17 22.02 26.46 10.66 46.13 25.91
LLaMAX Llama 3 8B 49.70 22.91 4.93 0.49 4.87 0.52
LLaMAX Llama 3 8B Alpaca 60.91 20.48 28.94 12.44 47.43 26.91
EMMA-500 Llama 3 8B Mono 62.34 22.16 41.29 22.03 53.16 32.55
EMMA-500 Llama 3 8B Bi 39.65 26.72 45.22 25.99 56.69 36.72
EMMA-500 Llama 3.1 8B Mono 26.17 19.62 40.32 21.06 50.34 29.51
EMMA-500 Llama 3.1 8B Bi 63.33 25.42 45.07 25.80 55.84 35.91
3.3 Model Adaptability
It is unrealistic to expect a single model to perform the best across all tasks and benchmarks, given the inherent trade-offs in multilingual generalization, task specialization, and resource distribution. In practice, we experience some performance drop of our models on certain benchmarks. In this section, we analyze model adaptability by examining how CPT impacts performance when applied to different base models across a wide range of languages. To quantify the effect, we compute the performance gain as the difference between the CPT model and its corresponding base model on each benchmark.
We compare LLaMA 2, LLaMA 3, and LLaMA 3.1 as base models for continual pre-training, each offering progressively larger and more recent training corpora. LLaMA 2 was trained on 2T tokens, while LLaMA 3 significantly expands coverage with over 15T tokens. LLaMA 3.1 further updates the model with extensive data and long-context fine-tuning. This progression allows us to assess the adaptability of CPT across different model generations and data scales. Figure˜4 shows the performance difference and the number of benchmarks where CPT models experience degradation.
CPT on Llama 2 observes a very small drop of BERTScore on MassiveSumm, where the score of the long subset decreases from 63.89 to 63.80 and the short one from 65.35 to 65.14. However, for Llama 3 and 3.1, both CPT on monolingual and bilingual mixes observe more cases of performance drops. For high-resource languages, Llama 3.1 degrades across more benchmarks than Llama 3. However, this trend diminishes notably when evaluating low-resource languages, where CPT remains effective to some extent.
This comparison provides insight into the effectiveness of continual pre-training with monolingual and bilingual data mixes in adapting different LLMs with various training to diverse linguistic settings. Our results align with the findings from Springer et al. (83) that over-trained language models are harder to fine-tune. At a massively multilingual scale, we corroborate that continuing pre-training on well-trained models—particularly those already optimized for high-resource languages, e.g., Llama 3 and 3.1—poses significant challenges when extending to hundreds of additional languages.
Figure 4: Model adaptability measured by the number of benchmarks on which CPT models are worse than the base model. CPT on LLaMA 2 (L2 Mono) shows a negligible BERTScore drop on MassiveSumm. More highly optimized models such as LLaMA 3 and 3.1 present greater challenges for effective continual pre-training compared to LLaMA 2, especially for high-resource languages, while the situation slightly eases for low-resource languages.
3.4 Per-Language Performance
Table 4: Percentage of top-$k$ rankings across four multilingual benchmarks. The darker color indicates the better. EMMA-500—especially with bilingual training—consistently outperforms baselines on Taxi1500 and Flores200. While Llama 3 and LlaMAX models yield higher accuracy on BELEBELE, EMMA-500 models cover more languages with improved relative performance.
BELEBELE SIB-200 Taxi1500 Flores200
Models Top 3 Top 5 Top 10 Top 3 Top 5 Top 10 Top 3 Top 5 Top 10 Top 3 Top 5 Top 10
Llama 3 8B 0.82 0.82 25.41 43.41 65.37 95.61 7.17 13.80 38.16 0.00 0.00 25.12
LLaMAX Llama 3 8B 0.00 1.64 11.48 0.49 7.32 35.12 6.10 19.38 58.73 0.00 0.00 0.99
LLaMAX Llama 3 8B Alpaca 0.00 1.64 20.49 5.37 23.90 77.07 3.45 6.10 10.62 0.00 0.00 44.83
EMMA-500 Llama 3 8B Mono 2.46 9.02 37.70 30.24 43.41 76.10 14.60 22.50 44.79 1.97 77.34 96.06
EMMA-500 Llama 3 8B Bi 4.10 13.93 44.26 2.93 10.24 20.49 32.45 47.38 78.23 88.18 98.52 100.00
Despite notable improvements, multilingual adaptation continues to present challenges. To scrutinize per-language performance, we evaluate each model by (1) counting the number of baseline models it outperforms across languages (from a total of 32 baselines; see Section˜D.3), and (2) calculating the percentage of languages where the model ranks among the top $k$ performers. This provides a fine-grained view of cross-lingual competitiveness.
Table˜4 presents the top-$k$ performance percentage across four multilingual benchmarks—BELEBELE, SIB-200, Taxi1500, and FLORES-200. The results demonstrate that EMMA-500 models, particularly those trained with bilingual data, achieve higher top-$k$ percentages on low-resource-heavy benchmarks like Taxi-1500 and FLORES-200. In contrast, base LLaMA 3 and LLaMAX variants exhibit more limited top-tier rankings. Notably, for BELEBELE, a machine comprehension benchmark, Llama 3 and 3.1 are not very strong models on this benchmark as evaluated in Section˜E.6. Existing CPT models, like LlaMAX Llama 3, and our CPT models obtained degraded performance. Despite having a decreased average accuracy, our models have better performance for more languages than the base model and the LlaMAX Llama 3 model.
We further look at the evaluation results of selected languages on BELEBELE in Table˜5 for a mix of low-resource (e.g., Khalkha Mongolian, Northern Sotho, Plateau Malagasy) and high-resource (e.g., English, French, Italian) languages. Table˜5 reports absolute scores and the delta () between each CPT model and the Llama 3 8B base model. Results show that while CPT variants generally improve scores for low-resource languages—especially in the EMMA-500 setups—they tend to suffer from notable performance degradation in high-resource languages. However, this trade-off is arguably acceptable as the main focus of this paper is to adapt LLMs into low-resource languages.
These analyses provide a deeper understanding of the strengths and limitations of massively multilingual adaptation and guide future improvements in multilingual language model training and evaluation. First, they show averaging scores across languages to compare model performance, while convenient, has several important limitations, e.g., obscuring disparities in performance across languages and resources, and ignoring per-language variance. Second, this comparison of per-language performance highlights the benefits of our CPT models in promoting cross-lingual competitiveness.
Table 5: Performance of selected languages in the BELEBELE benchmark. denotes the difference between the CPT model’s score and that of the base model. All models are based on Llama 3 8B. CPT on Llama 3 exhibits a significant drop in some high-resource languages, much bigger than the increase in some low-resource languages.
Model Average khk_Cyrl nso_Latn plt_Latn eng_Latn fra_Latn ita_Latn
Llama 3 8B 40.73 35.11 29.22 30.44 74.56 60.89 60.22
LLaMAX 36.96 34.67 30.33 34.11 62.33 48.78 47.33
-3.77 -0.44 1.11 3.67 -12.23 -12.11 -12.89
LLaMAX Alpaca 39.41 35.56 30.78 33.78 65.78 54.67 53.67
-1.32 0.45 1.56 3.34 -8.78 -6.22 -6.55
EMMA-500 Mono 39.73 38.44 37.00 34.78 56.00 51.11 49.56
-1.00 3.33 7.78 4.34 -18.56 -9.78 -10.66
EMMA-500 Bi 39.84 39.33 35.22 33.00 56.78 49.67 46.00
-0.89 4.22 6.00 2.56 -17.78 -11.22 -14.22
3.5 Overall Results
We evaluate the model on a comprehensive set of multilingual and bilingual benchmarks, assessing both language understanding and generation tasks. Tables˜6 and 7 present the results of deterministic and generation tasks. Our EMMA-500 models are the best at machine translation (Flores200) and competitive at text classification (Taxi1500 and SIB-200) and commonsense reasoning (XCOPA and XStoryCloze). While our models are competitive in many cases, we also observe some performance degradation, such asin math tasks Performance inevitably varies across tasks and languages, reflecting trade-offs in model capacity and data representation. Appendix˜E presents a deeper dive into the performance details across all tasks and languages evaluated, providing insights into model behavior across multilingual datasets and benchmarks. We also provide detailed per-language results, available at https://mala-lm.github.io/emma-500-v2
Table 6: Overall results of deterministic tasks including text classification, commonsense reasoning, natural language inference, reading comprehension, and math reasoning. XC: XCOPA; XSC: XStoryCloze; BELE: BELEBELE; ARC: the multilingual AI2 Reasoning Challenge; Dir.: MGSM by direct prompting; CoT: MGSM by CoT prompting. Underline and bold represents the absolute best, underline means the second best, and bold signifies the best within a specific group. Our EMMA-500 models are among the top 2 models on text classification and commonsense reasoning.
Model Classification Commonsense NLI Comprehension Math
SIB Taxi XC XSC XNLI BELE ARC Dir. CoT
Llama 2 7B 22.41 17.54 56.67 57.55 40.19 26.27 27.56 6.69 6.36
Llama 2 7B Chat 25.58 15.44 55.85 58.41 38.58 29.05 28.02 10.22 10.91
CodeLlama 2 7B 23.35 17.03 54.69 55.68 40.19 27.38 25.23 5.93 6.64
LLaMAX Llama 2 7B 10.61 23.52 54.38 60.36 44.27 23.09 26.09 3.35 3.62
LLaMAX Llama 2 7B Alpaca 27.89 15.09 56.60 63.83 45.09 24.48 31.06 5.05 6.35
MaLA-500 Llama 2 10B v1 23.25 25.27 53.09 53.07 38.11 22.96 21.16 0.91 0.73
MaLA-500 Llama 2 10B v2 19.30 23.39 53.09 53.07 38.11 22.96 21.16 0.91 0.73
YaYi Llama 2 7B 24.57 17.73 56.71 58.42 41.28 28.32 28.40 7.09 7.22
TowerBase Llama 2 7B 19.34 17.73 56.33 57.78 39.84 26.36 27.94 6.15 6.16
TowerInstruct Llama 2 7B 20.53 17.29 57.05 59.24 40.36 27.93 30.10 7.24 8.24
EMMA-500 Llama 2 7B 31.27 19.82 63.11 66.38 45.14 26.75 29.53 17.02 18.09
Occiglot Mistral 7B v0.1 32.69 22.26 56.67 58.10 42.35 30.16 29.77 13.31 14.07
Occiglot Mistral 7B v0.1 Instruct 34.31 18.76 56.55 59.39 40.81 32.05 30.88 22.76 22.16
BLOOM 7B 17.82 14.76 56.89 59.30 41.60 24.11 23.65 2.87 2.29
BLOOMZ 7B 29.73 16.96 54.87 57.12 37.13 39.32 23.95 2.55 2.15
YaYi 7B 35.76 16.12 56.64 60.67 39.87 37.97 24.44 2.76 3.02
Aya 23 8B 41.50 22.64 55.13 60.93 43.12 40.08 31.08 22.29 24.71
Aya Expanse 8B 57.01 18.73 56.38 64.80 45.56 46.98 36.56 43.02 41.45
Gemma 7B 58.21 13.83 63.64 65.01 42.58 43.37 38.68 38.22 35.78
Gemma 2 9B 46.25 18.05 66.33 67.67 46.74 54.49 44.15 32.95 44.69
Qwen 1.5 7B 47.95 7.29 59.44 59.85 39.47 41.83 28.93 31.56 30.36
Qwen 2 7B 54.95 21.87 60.31 61.46 42.77 49.31 33.82 48.95 51.47
Qwen 2.5 7B 53.89 17.87 61.71 62.06 43.31 54.11 35.30 53.78 55.60
Marco-LLM GLO 7B 64.15 21.99 62.45 63.87 43.99 53.95 36.34 51.85 52.02
Llama 3 8B 63.70 21.73 61.71 63.41 44.97 40.73 34.80 27.45 28.13
Llama 3.1 8B 61.42 20.20 61.71 63.58 45.62 45.19 34.93 28.36 27.31
LLaMAX Llama 3 8B 48.60 23.01 63.04 64.31 44.13 36.96 33.54 20.80 19.96
LLaMAX Llama 3 8B Alpaca 58.97 17.71 64.36 68.27 45.08 39.41 34.53 14.18 17.16
EMMA-500 Llama 3 8B Mono 60.62 22.32 66.20 67.36 44.15 39.73 33.22 23.53 25.33
EMMA-500 Llama 3 8B Bi 39.40 25.13 66.82 68.35 45.15 39.84 34.84 23.49 26.29
EMMA-500 Llama 3.1 8B Mono 26.16 19.71 65.38 67.64 39.98 38.86 34.00 24.95 27.35
EMMA-500 Llama 3.1 8B Bi 62.07 24.87 67.25 68.47 44.67 37.00 34.59 23.85 25.76
Table 7: Overall results of generation tasks including text summarization and machine translation. BERT: BERTScore; R-L: ROUGE-L. Underline and bold represents the absolute best, underline means the second best, and bold signifies the best within a specific group. Our EMMA-500 models trained with bilingual mix are the best-performing models on machine translation.
Model Flores200 MassiveSumm-L MassiveSumm-S XL-Sum
chrF++ BLEU chrF++ BLEU R-L BERT R-L BERT R-L BERT
Llama 2 7B 15.13 4.62 30.32 12.93 4.74 63.89 7.85 65.35 7.11 66.52
Llama 2 7B Chat 16.95 4.95 31.72 12.28 4.73 63.52 9.76 67.01 8.84 68.44
CodeLlama 2 7B 14.94 4.27 28.57 10.82 5.63 64.51 7.59 64.83 7.15 65.74
LLaMAX Llama 2 7B 7.42 0.80 13.66 1.99 4.56 62.69 5.22 63.06 5.29 64.59
LLaMAX Llama 2 7B Alpaca 28.35 12.51 42.27 22.29 4.61 62.76 10.71 67.92 10.11 69.24
MaLA-500 Llama 2 10B v1 6.08 0.60 13.60 2.29 4.39 64.50 4.97 63.51 5.45 63.96
MaLA-500 Llama 2 10B v2 6.38 0.54 15.44 2.87 4.37 64.66 5.02 63.75 5.44 64.28
YaYi Llama 2 7B 14.87 4.41 31.38 12.98 4.98 64.17 7.80 65.24 7.98 67.21
TowerBase Llama 2 7B 16.03 4.83 31.47 13.74 4.81 64.51 8.11 65.53 7.65 67.09
TowerInstruct Llama 2 7B 15.64 3.23 25.43 4.81 4.82 64.61 10.14 67.76 8.89 68.46
EMMA-500 Llama 2 7B 33.25 15.58 45.78 25.37 4.79 63.80 8.32 65.14 8.58 67.20
Occiglot Mistral 7B v0.1 16.10 4.32 31.13 13.12 5.14 63.95 8.16 63.65 7.33 66.20
Occiglot Mistral 7B v0.1 Instruct 15.80 3.99 31.65 11.61 5.16 63.50 7.82 63.79 8.31 66.96
BLOOM 7B 11.80 2.81 27.84 9.57 4.88 64.36 6.79 62.30 6.99 64.78
BLOOMZ 7B 16.10 7.44 34.74 20.22 2.91 57.20 3.28 29.75 11.15 69.82
YaYi 7B 13.50 4.37 21.36 4.82 4.95 64.24 8.28 65.44 12.06 69.74
Aya 23 8B 16.15 6.46 32.36 13.87 6.33 65.94 8.43 65.85 8.68 66.79
Aya Expanse 8B 23.89 6.88 36.86 13.12 7.44 67.66 9.24 67.68 10.51 68.73
Gemma 7B 23.05 9.05 43.68 23.79 6.18 62.14 8.35 62.25 6.70 64.52
Gemma 2 9B 26.48 12.09 38.87 23.15 5.86 59.70 7.86 58.11 7.38 65.45
Qwen 1.5 7B 17.77 5.87 35.87 15.58 6.09 59.19 8.49 62.70 9.58 69.13
Qwen 2 7B 17.17 5.56 37.61 17.39 6.65 56.31 8.63 56.14 10.18 69.35
Qwen 2.5 7B 17.49 5.72 38.89 18.95 7.44 61.62 9.04 58.91 10.41 69.69
Marco-LLM GLO 7B 23.34 9.27 44.57 25.17 6.57 57.15 8.10 57.45 11.46 70.41
Llama 3 8B 24.08 9.93 43.72 23.78 5.10 55.58 6.44 50.98 8.47 67.08
Llama 3.1 8B 24.69 10.11 44.10 24.19 5.41 56.09 6.77 54.96 8.57 66.97
LLaMAX Llama 3 8B 4.65 0.45 4.66 0.48 6.00 65.90 8.77 66.46 8.28 66.72
LLaMAX Llama 3 8B Alpaca 26.86 11.64 45.45 25.10 7.62 67.64 12.44 68.95 11.39 69.99
EMMA-500 Llama 3 8B Mono 38.34 20.33 50.86 30.38 5.38 61.06 7.18 63.46 9.11 66.12
EMMA-500 Llama 3 8B Bi 42.15 24.02 54.33 34.40 5.57 59.23 7.74 63.18 9.10 66.72
EMMA-500 Llama 3.1 8B Mono 37.41 19.44 48.12 27.57 5.44 61.89 7.21 63.36 9.66 67.21
EMMA-500 Llama 3.1 8B Bi 42.07 23.86 53.49 33.64 4.78 58.47 6.67 61.64 8.56 65.90
4 Related Work
Multilingual Continual pre-training
Multilingual language models have made significant progress in extending language understanding across diverse linguistic landscapes. Models such as mT5 (102) and BLOOM (78) have demonstrated strong multilingual capabilities by pre-training on extensive multilingual corpora. However, its performance on low-resource languages remains limited. Recent developments such as Qwen 2 (103) and Llama 3 (20) have further improved the efficiency and scalability of LLMs and shown stronger multilingual capabilities. Continual pre-training has emerged as an effective strategy for expanding language models’ linguistic and domain coverage without requiring full retraining. Prior works such as MaLA-500 (51), LlaMAX (54), and EMMA-500 (34) have demonstrated the feasibility of continual pre-training on diverse multilingual datasets. These approaches highlight the importance of incremental learning to incorporate new languages and improve cross-lingual transfer.
Despite these advancements, prior work has primarily focused on monolingual text pre-training, with limited emphasis on bilingual adaptation for improved cross-lingual transfer. Our work extends this line of research by introducing a bilingual translation corpus for continual pre-training and adapting Llama 3 and 3.1 (8B) to enhance bilingual and multilingual generalization, particularly for underrepresented languages.
Training on Bilingual Translation Data
Parallel texts have been used for pre-training LLMs. PolyLM (98) trained on 1 billion parallel multilingual data (0.16% of pre-training corpora). Poro (55) trained on 8 billion English-Finnish cross-lingual texts (slightly under 1% of pre-training corpora). Li et al. (48) trained small language models, i.e., BERT (19), GPT-2 (75), and BART (45), in a controlled setting with multilingual and bilingual texts using different learning objectives. In continual pre-training or fine-tuning, bilingual texts are also widely used. Ji et al. (35) showed continual pre-training of multilingual BART (86) with machine translation failed enhance cross-lingual representation learning. Xu et al. (101) fine-tuned LLMs with monolingual data and parallel data for machine translation. Kondo et al. (41) showed that continual pre-training and supervised fine-tuning on parallel data can enhance the accuracy in English-Japanese translation. Similarly, Fujii et al. (23) investigated cross-lingual continual pre-training on English and Japanese and showed that the use of parallel corpora enhanced translation performance. Li et al. (49) studied the data mixing strategies for continual pre-training using monolingual and parallel texts and code data across 30+ languages. While previous studies have explored bilingual and multilingual pre-training with parallel texts, most efforts have been limited to a small subset of languages or specific translation tasks. Unlike PolyLM and Poro, which incorporate a modest fraction of bilingual data into predominantly monolingual training, our work systematically integrates bilingual texts across 500+ languages.
5 Conclusion
This work advances massively multilingual adaptation of LLMs using bilingual translation data across 2,500+ language pairs (500+ languages). The four released EMMA-500 models using Llama 3 and 3.1 trained on both monolingual and bilingual data mixes, including the newly compiled MaLA translation corpus, establish new benchmarks in multilingual coverage while maintaining competitive performance across 7 diverse tasks. Notably, they achieve state-of-the-art results on machine translation while showing robust generalization to text classification and reasoning tasks. Despite advancements, achieving consistently high performance across diverse benchmarks is constrained by linguistic and task variability, which highlights intrinsic tensions between scale and specialization in multilingual NLP.
Limitations
Multilingual Benchmark
Multilingual language models are designed to accommodate users across various linguistic and cultural backgrounds. However, many widely used multilingual evaluation benchmarks, including some in this research, are developed through human or machine translation. As a result, they often emphasize knowledge and subject matter primarily from English-speaking sources, with potential translation-related distortions that can undermine the accuracy of model assessments (11). This discrepancy highlights the need for more comprehensive, native-language test sets that better reflect the full range of linguistic diversity. We encourage collaborative efforts to create large-scale benchmarks that offer a more reliable evaluation of these models across different languages and cultures.
Human Evaluation
Although human assessment is a valuable tool, it presents challenges such as variability, subjectivity, and high costs—especially when evaluating models across numerous languages. Recruiting expert annotators proficient in less common languages is particularly difficult, making large-scale human evaluation unfeasible. Even assessing a limited selection of languages requires significant resources. While this study acknowledges these constraints, we recognize that human evaluations play a crucial role in supplementing automated assessment methods. Given these limitations, we rely on automatic evaluation tools to ensure scalability and consistency, despite their imperfections.
Model Performance
Despite its strengths in multilingual processing, our models face challenges in areas such as mathematical reasoning and machine reading comprehension. Its performance on machine-translated mathematical benchmarks remains limited, likely due to the inherent difficulty of numerical reasoning and translation artifacts that may obscure problem clarity. Similarly, while our model achieves improvements in various NLP tasks, it struggles with reading comprehension, which demands deep contextual understanding and logical inference. Addressing these weaknesses will require further refinements, such as incorporating domain-specific training data or exploring alternative model architectures optimized for these challenges. Besides, averaging scores across languages can mask important disparities, such as strong performance on high-resource languages and poor results on low-resource ones. It also ignores linguistic diversity and per-language difficulty, potentially leading to misleading conclusions about overall model performance.
Real-world Usage
This research primarily focuses on enhancing continual training with bilingual texts and improving language model performance through continual pre-training. However, the model is not yet suitable for deployment in real-world applications. It has not undergone thorough human alignment processes or adversarial robustness testing (red-teaming) to ensure safety and reliability. While our work contributes to advancements in multilingual NLP, additional refinements—such as aligning the model with human preferences and conducting rigorous safety evaluations—are necessary before it can be practically implemented.
Data Mix
We examine two manually constructed data mixes to assess the impact of CPT with bilingual translation data. While numerous alternative configurations are possible, systematically validating the effectiveness of different data mixes or searching for the optimal data mixes requires a lot of computing resources, which are not affordable for a small research team like us. Nonetheless, the two data mixes carefully decided by us already demonstrate measurable improvements in the downstream evaluation, which highlight the utility of our design choices.
Model Training
Model training involves a few hyperparameters, and a wide range of base models can be selected for continual pre-training. When preparing the bilingual translation data for training, the number of lines is intuitively chosen. However, an exhaustive search over all possible configuration combinations—including data volume, model choice, and hyperparameter settings—would incur prohibitive computational costs. Consequently, we focus on a limited set of configurations that are feasible under available resources.
Community Collaboration
This study was conducted without direct community collaboration. Nonetheless, we recognize the value of community collaboration and are open to future partnerships with researchers and practitioners in this area.
Acknowledgment
The work has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No 101070350 and from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee [grant number 10052546], and the Digital Europe Programme under grant agreement No 101195233. The authors wish to acknowledge CSC – IT Center for Science, Finland, for computational resources. We thank Indraneil Paul for his contributions to preparing the code data.
Appendix A Frequently Asked Questions
Goals and Achievements
Q: What is the main goal of this paper?
The main goal of this paper, as suggested by its title, is the perform massively multilingual adaptation of large language models using bilingual translation data. Specifically, we compile a massive parallel corpus, compose two data mixes, perform continual pre-training of large language models and comprehensive evaluation, and examine interesting research questions such as the effect of using bilingual translation data.
Q: What are the main achievements and the scientific novelty?
A: Our main achievements are summarized as the contributions as written in the Section˜1 Introduction. For continual pre-training, we do not modify the model architecture—we do not claim this novelty. However, our scientific novelty is more on the resource construction side, including the creation of MaLA translation corpus and continually pre-trained EMMA-500 models based on Llama 3 and 3.1 on two diverse data mixes. Besides, our empirical evaluation reveals some valuable findings that can potentially promote future development.
Data
Q: How was the data mix tuned?
Data mix optimization is a distinct research area. Our focus was on creating a robust bilingual translation corpus through careful curation for language and content diversity. While we didn’t perform specific data mix ablations, we ensured balanced representation across languages and domains and compared two data mixes with or without bilingual translation data to examine the effect of including parallel data for continual pre-training. Searching for the optimal data composition requires a lot of computing resources and remains an interesting direction for future work.
Q: Does whitespace cleaning only apply to whitespace-tokenized languages, or do other languages require different cleaning methods?
A: Whitespace cleaning is mainly relevant for whitespace-tokenized languages (like English) to standardize text for processing. However, non-whitespace-tokenized languages (e.g., Chinese, Japanese, Korean) may not need this step but often require other pre-processing, such as character normalization or segmentation.
Q: How were the existing datasets and their languages selected for inclusion?
A: The datasets and languages were chosen based on Ji et al. (34) adopting the the following criteria:
  • Language Coverage: Emphasis on broad multilingual representation, including both high- and low-resource languages.
  • Data Quality: Preference for high-quality, diverse content (books, papers, code) to enhance model generalization.
  • Task Relevance: Alignment with goals for continual pre-training and multilingual adaptation (e.g., translation, classification, reasoning).
  • Availability: Prioritization of publicly accessible datasets to ensure transparency and reproducibility.
Evaluation
Q: How do we determine the number of shots for in-context learning?
A: We use the same number of shots as Ji et al. (34), where the number of shots used in evaluation follows the setup from Llama 2, which demonstrates strong performance on high-resource languages with this configuration. Rather than performing an exhaustive grid search for optimal shot count, we maintain this established approach. We provide the model weights, enabling researchers to experiment with different numbers of shots according to their specific evaluation needs.
Q: Why no human evaluation?
A: Human evaluation faces scalability challenges in multilingual settings due to subjectivity, cultural biases, and high costs—especially for low-resource languages requiring specialized annotators. While valuable, these limitations make automated metrics more practical for our large-scale study. We acknowledge human evaluation’s importance, but emphasize its unsustainability for massively multilingual work.
Q: How reliable is machine-generated benchmark?
A: Machine-generated benchmarks provide scalable, consistent, and objective evaluation across many tasks and languages, enabling efficient model comparison. However, they often lack nuance in assessing creativity, contextual appropriateness, and cultural sensitivity. Potential biases in generation or evaluation metrics may also affect reliability. The study on the reliability of machine-generated benchmarks is out of the scope of this paper.
Q: Why no safety-related evaluation?
A: Safety evaluation is a distinct research area. Our work focuses on CPT, releasing a multilingual base model (not instruction-tuned) and corpus. As this isn’t an end-user product, safety wasn’t a primary focus. Existing benchmarks like Wang et al. (96) and Aakanksha et al. (1) cover a few languages and aren’t ideal for massively multilingual settings.
Comparison with Baselines
Q: Is comparing models on unsupported languages unfair?
A: We evaluate all models across diverse languages, including unsupported ones, to assess generalization capabilities. There’s no established standard for “fair” baselines in such cross-lingual comparisons. Evaluating on unsupported languages stresses the model’s zero-shot transferability.
Q: What is the relation to the EMMA-500 Llama 2 model (34)?
A: This work is a follow-up of the EMMA-500 Llama 2 model. Our models follow most of the settings of the EMMA-500 Llama 2 model, but use additional bilingual translation data, continue pre-pretraining on Llama 3 and 3.1 models, and study on two big data mixes. Moreover, our analyses reveals valuable findings.
Q: Why not comparing with model XXX?
A: Our model selection prioritizes representative architectures that have diverse capabilities. Particularly, we focus on open-weight decoder-only models from 7-9B parameters. Our current exclusion reflects either the overlap with existing benchmarks or technical constraints. Also, some newer models would come up. We actively welcome feedback to expand coverage of baselines.
Others
Q: Will the corpus, model weights, benchmarks, and scripts be publicly available?
A: Yes. We release the MaLA translation corpus, all model weights, and code. Benchmarks and processing scripts are open-sourced by their original authors.
Q: How do we categorize languages into resource groups?
A: We follow the resource groups used by Ji et al. (34), where languages are grouped into five categories (high to low resource) based on token counts in the MaLA monolingual corpus. This data-driven approach reflects actual text availability. We include ISO codes and writing systems in the MaLA parallel corpus for standardization across NLP applications.
Q: Why don’t we separate metrics for officially supported vs. unseen languages?
A: While possible for models declaring supported languages, this can’t be done for LLaMA or GEMMA which lack such specifications. We include per-language results in the Appendix and will share all model generations on our project website (https://mala-lm.github.io/emma-500-v2), enabling community analysis with custom metrics while keeping the paper concise.
Appendix B Details of MaLA Translation Corpus
B.1 Data Sources
Table˜8 lists the corpora and collections we use as bilingual data sources in this work. A data source being bilingual means that it contains corresponding, or parallel, text records in two languages. This usually means one side has been translated, either by a machine or a human. Pedantically, massively parallel bilingual corpora can be considered a special case of multilingual corpora containing sentences in different languages that share the same meaning. In this work, some data sources have parallel text data in more than two languages, but in such cases, we always iteratively determine one as the source and one as the target language, i.e., create bilingual corpora out of them. In some source datasets, the source language is not explicitly stated, and as such we pick one language to regard as the source language without actual knowledge of the direction of translation. In such cases, we tend to pick a high-resource language, such as English, as the source language.
Metadata
In the case of bilingual corpora, we define each record in the JSONL file to consist of the fields src_lang, src_txt, tgt_lang, tgt_txt, url, collection, source, original_src_lang, and original_tgt_lang. The contents of these fields are as follows. The fields src_txt and tgt_txt contain the parallel text data in the source and target language, respectively, and src_lang and tgt_lang are the ISO 639-3 language codes of those languages. Identically with the monolingual corpora, url contains a URL indicating the web address from which the text data has been extracted, if available, collection the name of the collection, and source the name of a more specific part of the collection which the text data was extracted from. Lastly, original_src_lang and original_tgt_lang contain the language denotations of the source and target language, respectively, as they are given in the source data.
Table 8: Source datasets used for compiling the MaLA bilingual translation data.
Dataset Languages URL
CMU Haitian Creole (93) Haitian Creole http://www.speech.cs.cmu.edu/haitian/
English-Luganda Parallel Corpus (63) Luganda-English https://zenodo.org/records/4764039
Ewe language corpus (27) Ewe—English https://www.kaggle.com/datasets/yvicherita/ewe-language-corpus
Nunavut Hansard Inuktitut–English Parallel Corpus 3.0 (37) Inuktitut–English https://nrc-digital-repository.canada.ca/eng/view/object/?id=c7e34fa7-7629-43c2-bd6d-19b32bf64f60
AmericasNLP 2021 (28, 95, 25) 10 https://turing.iimas.unam.mx/americasnlp/americasnlp_2021.html;https://github.com/AmericasNLP/americasnlp2021
AmericasNLP 2023 (57, 28, 12, 65, 95) 11 https://turing.iimas.unam.mx/americasnlp/2023_st.html
AmericasNLP 2022 4 https://github.com/AmericasNLP/americasnlp2022
Indigenous Languages Corpora (21) Cree https://github.com/EdTeKLA/IndigenousLanguages_Corpora
ACES (5) 146 https://huggingface.co/datasets/nikitam/ACES
ChrEn (105) 2 https://huggingface.co/datasets/chr_en
NusaX-MT (99) 12 https://huggingface.co/datasets/indonlp/NusaX-MT
Lego-MT (104) 433 https://github.com/CONE-MT/Lego-MT
Tatoeba (90) 487 https://github.com/Helsinki-NLP/Tatoeba-Challenge/tree/master/data
NLLB (64) 202 https://huggingface.co/datasets/allenai/nllb
Lacuna Project (59) 1 https://github.com/masakhane-io/lacuna_pos_ner
lafand-mt (2) 21 https://github.com/masakhane-io/lafand-mt/tree/main
B.2 Supported Languages
Table˜9 shows the language codes of the MaLA corpus (34), where “unseen” means the languages are not used for training EMMA-500. The classification system for token counts categorizes language resources based on their size into five distinct tiers: “high” for resources exceeding 1 billion tokens, indicating a vast amount of data; “medium-high” for those with more than 100 million tokens, reflecting a substantial dataset; “medium” for resources that contain over 10 million tokens, representing a moderate size; “medium-low” for datasets with over “1 million tokens”, indicating a smaller yet significant amount of data; and finally, “low” for resources containing less than 1 million tokens, which suggests a minimal data presence. This hierarchy helps in understanding the scale and potential utility of the language resources available.
Table 9: Languages by resource groups categorized by counting the number of tokens in the MaLA monolingual corpus (34). “Unseen” means those languages are not used for continual pre-training in this paper.
Category Languages Language Codes
high 27 fra_Latn, mon_Cyrl, kat_Geor, tgk_Cyrl, kaz_Cyrl, glg_Latn, hbs_Latn, kan_Knda, mal_Mlym, rus_Cyrl, cat_Latn, hye_Armn, guj_Gujr, slv_Latn, fil_Latn, bel_Cyrl, isl_Latn, nep_Deva, mlt_Latn, pan_Guru, afr_Latn, urd_Arab, mkd_Cyrl, aze_Latn, deu_Latn, eng_Latn, ind_Latn
low 210 prs_Arab, nqo_Nkoo, emp_Latn, pfl_Latn, teo_Latn, gpe_Latn, izz_Latn, shn_Mymr, hak_Latn, pls_Latn, evn_Cyrl, djk_Latn, toj_Latn, nog_Cyrl, ctu_Latn, tca_Latn, jiv_Latn, ach_Latn, mrj_Latn, ajp_Arab, apc_Arab, tab_Cyrl, hvn_Latn, tls_Latn, bak_Latn, ndc_Latn, trv_Latn, top_Latn, kjh_Cyrl, guh_Latn, mni_Mtei, csy_Latn, noa_Latn, dov_Latn, bho_Deva, kon_Latn, hne_Deva, kcg_Latn, mni_Beng, hus_Latn, pau_Latn, jbo_Latn, dtp_Latn, kmb_Latn, hau_Arab, pdc_Latn, nch_Latn, acf_Latn, bim_Latn, ixl_Latn, dty_Deva, kas_Arab, lrc_Arab, alz_Latn, lez_Cyrl, lld_Latn, tdt_Latn, acm_Arab, bih_Deva, mzh_Latn, guw_Latn, rop_Latn, rwo_Latn, ahk_Latn, qub_Latn, kri_Latn, gub_Latn, laj_Latn, sxn_Latn, luo_Latn, tly_Latn, pwn_Latn, mag_Deva, xav_Latn, bum_Latn, ubu_Latn, roa_Latn, mah_Latn, tsg_Latn, gcr_Latn, arn_Latn, csb_Latn, guc_Latn, bat_Latn, knj_Latn, cre_Latn, bus_Latn, anp_Deva, aln_Latn, nah_Latn, zai_Latn, kpv_Cyrl, enq_Latn, gvl_Latn, wal_Latn, fiu_Latn, swh_Latn, crh_Latn, nia_Latn, bqc_Latn, map_Latn, atj_Latn, npi_Deva, bru_Latn, din_Latn, pis_Latn, gur_Latn, cuk_Latn, zne_Latn, cdo_Latn, lhu_Latn, pcd_Latn, mas_Latn, bis_Latn, ncj_Latn, ibb_Latn, tay_Latn, bts_Latn, tzj_Latn, bzj_Latn, cce_Latn, jvn_Latn, ndo_Latn, rug_Latn, koi_Cyrl, mco_Latn, fat_Latn, olo_Latn, inb_Latn, mkn_Latn, qvi_Latn, mak_Latn, ktu_Latn, nrm_Latn, kua_Latn, san_Latn, nbl_Latn, kik_Latn, dyu_Latn, sgs_Latn, msm_Latn, mnw_Latn, zha_Latn, sja_Latn, xal_Cyrl, rmc_Latn, ami_Latn, sda_Latn, tdx_Latn, yap_Latn, tzh_Latn, sus_Latn, ikk_Latn, bas_Latn, nde_Latn, dsb_Latn, seh_Latn, knv_Latn, amu_Latn, dwr_Latn, iku_Cans, uig_Latn, bxr_Cyrl, tcy_Knda, mau_Latn, aoj_Latn, gor_Latn, cha_Latn, fip_Latn, chr_Cher, mdf_Cyrl, arb_Arab, quw_Latn, shp_Latn, spp_Latn, frp_Latn, ape_Latn, cbk_Latn, mnw_Mymr, mfe_Latn, jam_Latn, lad_Latn, awa_Deva, mad_Latn, ote_Latn, shi_Latn, btx_Latn, maz_Latn, ppk_Latn, smn_Latn, twu_Latn, blk_Mymr, msi_Latn, naq_Latn, tly_Arab, wuu_Hani, mos_Latn, cab_Latn, zlm_Latn, gag_Latn, suz_Deva, ksw_Mymr, gug_Latn, nij_Latn, nov_Latn, srm_Latn, jac_Latn, nyu_Latn, yom_Latn, gui_Latn
medium 68 tha_Thai, kat_Latn, lim_Latn, tgk_Arab, che_Cyrl, lav_Latn, xho_Latn, war_Latn, nan_Latn, grc_Grek, orm_Latn, zsm_Latn, cnh_Latn, yor_Latn, arg_Latn, tgk_Latn, azj_Latn, tel_Latn, slk_Latn, pap_Latn, zho_Hani, sme_Latn, tgl_Latn, uzn_Cyrl, als_Latn, san_Deva, azb_Arab, ory_Orya, lmo_Latn, bre_Latn, mvf_Mong, fao_Latn, oci_Latn, sah_Cyrl, sco_Latn, tuk_Latn, aze_Arab, hin_Deva, haw_Latn, glk_Arab, oss_Cyrl, lug_Latn, tet_Latn, tsn_Latn, hrv_Latn, gsw_Latn, arz_Arab, vec_Latn, mon_Latn, ilo_Latn, ctd_Latn, ben_Beng, roh_Latn, kal_Latn, asm_Beng, srp_Latn, bod_Tibt, hif_Latn, rus_Latn, nds_Latn, lus_Latn, ido_Latn, lao_Laoo, tir_Ethi, chv_Cyrl, wln_Latn, kaa_Latn, pnb_Arab
medium-high 79 div_Thaa, som_Latn, jpn_Japn, hat_Latn, sna_Latn, heb_Hebr, bak_Cyrl, nld_Latn, tel_Telu, kin_Latn, msa_Latn, gla_Latn, bos_Latn, dan_Latn, smo_Latn, ita_Latn, mar_Deva, pus_Arab, srp_Cyrl, spa_Latn, lat_Latn, hmn_Latn, sin_Sinh, zul_Latn, bul_Cyrl, amh_Ethi, ron_Latn, tam_Taml, khm_Khmr, nno_Latn, cos_Latn, fin_Latn, ori_Orya, uig_Arab, hbs_Cyrl, gle_Latn, cym_Latn, vie_Latn, kor_Hang, lit_Latn, yid_Hebr, ara_Arab, sqi_Latn, pol_Latn, tur_Latn, swa_Latn, hau_Latn, ceb_Latn, eus_Latn, kir_Cyrl, mlg_Latn, jav_Latn, snd_Arab, sot_Latn, por_Latn, uzb_Cyrl, fas_Arab, nor_Latn, est_Latn, hun_Latn, ibo_Latn, ltz_Latn, swe_Latn, tat_Cyrl, ast_Latn, mya_Mymr, uzb_Latn, sun_Latn, ell_Grek, ces_Latn, mri_Latn, ckb_Arab, kur_Latn, kaa_Cyrl, nob_Latn, ukr_Cyrl, fry_Latn, epo_Latn, nya_Latn
medium-low 162 aym_Latn, rue_Cyrl, rom_Latn, dzo_Tibt, poh_Latn, sat_Olck, ary_Arab, fur_Latn, mbt_Latn, bpy_Beng, iso_Latn, pon_Latn, glv_Latn, new_Deva, gym_Latn, bgp_Latn, kac_Latn, abt_Latn, quc_Latn, otq_Latn, sag_Latn, cak_Latn, avk_Latn, pam_Latn, meo_Latn, tum_Latn, bam_Latn, kha_Latn, syr_Syrc, kom_Cyrl, nhe_Latn, bal_Arab, srd_Latn, krc_Cyrl, lfn_Latn, bar_Latn, rcf_Latn, nav_Latn, nnb_Latn, sdh_Arab, aka_Latn, bew_Cyrl, bbc_Latn, meu_Latn, zza_Latn, ext_Latn, yue_Hani, ekk_Latn, xmf_Geor, nap_Latn, mzn_Arab, pcm_Latn, lij_Latn, myv_Cyrl, scn_Latn, dag_Latn, ban_Latn, twi_Latn, udm_Cyrl, som_Arab, nso_Latn, pck_Latn, crs_Latn, acr_Latn, tat_Latn, afb_Arab, uzs_Arab, hil_Latn, mgh_Latn, tpi_Latn, ady_Cyrl, pag_Latn, kiu_Latn, ber_Latn, iba_Latn, ksh_Latn, plt_Latn, lin_Latn, chk_Latn, tzo_Latn, tlh_Latn, ile_Latn, lub_Latn, hui_Latn, min_Latn, bjn_Latn, szl_Latn, kbp_Latn, inh_Cyrl, que_Latn, ven_Latn, vls_Latn, kbd_Cyrl, run_Latn, wol_Latn, ace_Latn, ada_Latn, kek_Latn, yua_Latn, tbz_Latn, gom_Latn, ful_Latn, mrj_Cyrl, abk_Cyrl, tuc_Latn, stq_Latn, mwl_Latn, tvl_Latn, quh_Latn, gom_Deva, mhr_Cyrl, fij_Latn, grn_Latn, zap_Latn, mam_Latn, mps_Latn, tiv_Latn, ksd_Latn, ton_Latn, bik_Latn, vol_Latn, ava_Cyrl, tso_Latn, szy_Latn, ngu_Latn, hyw_Armn, fon_Latn, skr_Arab, kos_Latn, tyz_Latn, kur_Arab, srn_Latn, tyv_Cyrl, bci_Latn, vep_Latn, crh_Cyrl, kpg_Latn, hsb_Latn, ssw_Latn, zea_Latn, ewe_Latn, ium_Latn, diq_Latn, ltg_Latn, nzi_Latn, guj_Deva, ina_Latn, pms_Latn, bua_Cyrl, lvs_Latn, eml_Latn, hmo_Latn, kum_Cyrl, kab_Latn, chm_Cyrl, cor_Latn, cfm_Latn, alt_Cyrl, bcl_Latn, ang_Latn, frr_Latn, mai_Deva
unseen 393 rap_Latn, pmf_Latn, lsi_Latn, dje_Latn, bkx_Latn, ipk_Latn, syw_Deva, ann_Latn, bag_Latn, bat_Cyrl, chu_Cyrl, gwc_Arab, adh_Latn, szy_Hani, shi_Arab, njy_Latn, pdu_Latn, buo_Latn, cuv_Latn, udg_Mlym, bax_Latn, tio_Latn, kjb_Latn, taj_Deva, lez_Latn, olo_Cyrl, rnl_Latn, bri_Latn, inh_Latn, kas_Cyrl, wni_Latn, anp_Latn, tsc_Latn, mgg_Latn, udi_Cyrl, mdf_Latn, agr_Latn, xty_Latn, llg_Latn, nge_Latn, gan_Latn, tuv_Latn, stk_Latn, nut_Latn, thy_Thai, lgr_Latn, hnj_Latn, dar_Cyrl, aia_Latn, lwl_Thai, tnl_Latn, tvs_Latn, jra_Khmr, tay_Hani, gal_Latn, ybi_Deva, snk_Arab, gag_Cyrl, tuk_Cyrl, trv_Hani, ydd_Hebr, kea_Latn, gbm_Deva, kwi_Latn, hro_Latn, rki_Latn, quy_Latn, tdg_Deva, zha_Hani, pcg_Mlym, tom_Latn, nsn_Latn, quf_Latn, jmx_Latn, kqr_Latn, mrn_Latn, bxa_Latn, abc_Latn, mve_Arab, lfa_Latn, qup_Latn, yin_Latn, roo_Latn, mrw_Latn, nxa_Latn, yrk_Cyrl, bem_Latn, kvt_Latn, csw_Cans, bjr_Latn, mgm_Latn, ngn_Latn, pib_Latn, quz_Latn, awb_Latn, myk_Latn, otq_Arab, ino_Latn, tkd_Latn, bef_Latn, bug_Bugi, aeu_Latn, nlv_Latn, dty_Latn, bkc_Latn, mmu_Latn, hak_Hani, sea_Latn, mlk_Latn, cbr_Latn, lmp_Latn, tnn_Latn, qvz_Latn, pbt_Arab, cjs_Cyrl, mlw_Latn, mnf_Latn, bfm_Latn, dig_Latn, thk_Latn, zxx_Latn, lkb_Latn, chr_Latn, pnt_Latn, vif_Latn, fli_Latn, got_Latn, hbb_Latn, tll_Latn, bug_Latn, kxp_Arab, qaa_Latn, krr_Khmr, kjg_Laoo, isu_Latn, kmu_Latn, gof_Latn, sdk_Latn, mne_Latn, baw_Latn, idt_Latn, xkg_Latn, mgo_Latn, dtr_Latn, kms_Latn, ffm_Latn, hna_Latn, nxl_Latn, bfd_Latn, odk_Arab, miq_Latn, mhx_Latn, kam_Latn, yao_Latn, pnt_Grek, kby_Latn, kpv_Latn, kbx_Latn, cim_Latn, qvo_Latn, pih_Latn, nog_Latn, nco_Latn, rmy_Cyrl, clo_Latn, dmg_Latn, aaa_Latn, rel_Latn, ben_Latn, loh_Latn, thl_Deva, chd_Latn, cni_Latn, cjs_Latn, lbe_Latn, ybh_Deva, zxx_Zyyy, awa_Latn, gou_Latn, xmm_Latn, nqo_Latn, rut_Cyrl, kbq_Latn, tkr_Latn, dwr_Ethi, ckt_Cyrl, ady_Latn, yea_Mlym, nhx_Latn, niv_Cyrl, bwt_Latn, xmg_Latn, chy_Latn, mfj_Latn, hre_Latn, bbk_Latn, shn_Latn, lrc_Latn, qvc_Latn, muv_Mlym, mdr_Latn, luy_Latn, lzh_Hani, fuh_Latn, mle_Latn, brx_Deva, pex_Latn, kau_Latn, yrk_Latn, hin_Latn, ekm_Latn, msb_Latn, unr_Orya, cac_Latn, chp_Cans, ckt_Latn, bss_Latn, lts_Latn, bbj_Latn, ttt_Cyrl, kwu_Latn, smn_Cyrl, kpy_Cyrl, tod_Latn, wbm_Latn, tcy_Latn, arc_Syrc, nst_Latn, tuz_Latn, bob_Latn, bfn_Latn, pli_Deva, snl_Latn, kwd_Latn, lgg_Latn, nza_Latn, wbr_Deva, lan_Latn, kmz_Latn, bzi_Thai, hao_Latn, nla_Latn, qxr_Latn, ken_Latn, tbj_Latn, blk_Latn, ybb_Latn, nwe_Latn, gan_Hani, snk_Latn, kak_Latn, tpl_Latn, hla_Latn, tks_Arab, pea_Latn, bya_Latn, enc_Latn, jgo_Latn, tnp_Latn, aph_Deva, bgf_Latn, brv_Laoo, nod_Thai, niq_Latn, nwi_Latn, xmd_Latn, gbj_Orya, syr_Latn, ify_Latn, xal_Latn, bra_Deva, cgc_Latn, bhs_Latn, pwg_Latn, ang_Runr, oki_Latn, qve_Latn, qvm_Latn, bkm_Latn, bkh_Latn, niv_Latn, zuh_Latn, mry_Latn, fiu_Cyrl, ssn_Latn, rki_Mymr, sox_Latn, yav_Latn, nyo_Latn, dag_Arab, qxh_Latn, bze_Latn, myx_Latn, zaw_Latn, ddg_Latn, wnk_Latn, bwx_Latn, mqy_Latn, lad_Hebr, boz_Latn, lue_Latn, ded_Latn, pli_Latn, avk_Cyrl, wms_Latn, sgd_Latn, azn_Latn, ajz_Latn, psp_Latn, jra_Latn, smt_Latn, ags_Latn, csw_Latn, wtk_Latn, emp_Cyrl, koi_Latn, tkr_Cyrl, amp_Latn, ymp_Latn, mfh_Latn, tdb_Deva, omw_Latn, khb_Talu, doi_Deva, gld_Cyrl, ava_Latn, chu_Latn, dnw_Latn, azo_Latn, dug_Latn, bce_Latn, kmr_Latn, kpy_Armn, abq_Cyrl, trp_Latn, ewo_Latn, the_Deva, hig_Latn, pkb_Latn, mxu_Latn, oji_Latn, tnt_Latn, mzm_Latn, mns_Cyrl, lbe_Cyrl, qvh_Latn, kmg_Latn, sps_Latn, brb_Khmr, tah_Latn, sxb_Latn, mkz_Latn, mgq_Latn, got_Goth, lns_Latn, arc_Latn, akb_Latn, skr_Latn, nsk_Cans, sml_Latn, pce_Mymr, eee_Thai, lhm_Deva, yux_Cyrl, bqm_Latn, bcc_Arab, nas_Latn, agq_Latn, xog_Latn, tsb_Latn, fub_Latn, mqj_Latn, nsk_Latn, bxr_Latn, dln_Latn, ozm_Latn, rmy_Latn, cre_Cans, kim_Cyrl, cuh_Latn, ngl_Latn, yas_Latn, bud_Latn, miy_Latn, ame_Latn, pnz_Latn, raj_Deva, enb_Latn, cmo_Khmr, saq_Latn, tpu_Khmr, eve_Cyrl, cdo_Hani
B.3 Token Counts
Figure˜5 shows the numbers of segments and tokens across all language pairs in the MaLA translation corpus.
(a) Number of tokens
(b) Number of segments
Figure 5: Numbers of segments and tokens across all language pairs in the MaLA bilingual translation corpus.
Appendix C Details of Data Mixes
Table˜10 provides how different data types and resource categories contribute to the overall dataset composition. Our data mixes aim to make a balanced distribution over high- and low-resources. However, due to the nature of high-resource languages, they still contribute to a large portion. For example, English research papers form a substantial portion. Medium-high and medium-resource monolingual and bilingual texts contribute a lot. Medium and low-resource languages consist of 33% and 19% of bilingual and monolingual mixes, respectively.
C.1 Additional Code Data
Code Data
We source code data from a de-duplicated version of The Stack (40) and oversample files from the algorithmic coding (73) and data science domains  [^11] https://huggingface.co/datasets/AlgorithmicResearchGroup/arxiv_research_code owing to the benefits of training models on self-contained code (24). We discard all non-data programming languages that occur fewer than 50k times, with the exception of llvm, following prior work detailing its importance in multi-lingual and low-resource code generation (85, 67). We also discard samples that manifest code in rare but valid extensions. Finally, we source from data-heavy formats but follow precedent (53) and subsample them more aggressively. The surviving data is filtered as follows:
  • For files forked more than 25 times, we retain them if the average line length is less than 120, the maximum line length is less than 300, and the alphanumeric fraction is more than 30%.
  • For files forked between 15 and 25 times, we retain them if the average line length is less than 90, the maximum line length is less than 150, and the alphanumeric fraction is more than 40%.
  • For files forked less than 15 times, we retain them if the average line length is less than 80, the maximum line length is less than 120, and the alphanumeric fraction is more than 45%.
Code-Adjacent Procedural Data
We augment our data mix with procedural code-adjacent data ranging from instructive data such as StackOverflow QnAs, library documentation and tutorials  [^12] https://huggingface.co/datasets/BEE-spoke-data/code-tutorials-en to Jupyter notebooks, synthetic code textbooks (84), GitHub commits, and issue descriptions (53). We further mirror Paul et al. (68) in sourcing parallel math problem solutions to code data.
Table 10: Composition of training data mixtures showing original and final token counts by data type, with sampling rates and distribution percentages across bilingual and monolingual configurations.
Type Category Sample Token Counts Percentage of Mixes
Rate Original Final Bilingual Monolingual
instruction EN 0.1 9,204,199,807 920,419,981 0.32% 0.47%
high 0.2 39,403,448,029 7,880,689,606 2.72% 4.01%
medium-high 0.5 30,651,187,534 15,325,593,767 5.28% 7.81%
medium 5.0 1,444,764,863 7,223,824,315 2.49% 3.68%
medium-low 20.0 47,691,495 953,829,900 0.33% 0.49%
low 50.0 3,064,796 153,239,800 0.05% 0.08%
code/reasoning 1.0 612,208,775 612,208,775 0.21% 0.31%
code code 1.0 43,478,432,765 43,478,432,765 14.99% 22.15%
book Gutenberg 1.0 5,173,357,710 5,173,357,710 1.78% 2.64%
paper EN 1.0 38,256,934,181 38,256,934,181 13.19% 19.49%
ZH 1.0 61,787,372 61,787,372 0.02% 0.03%
monolingual EN 0.1 3,002,029,817 300,202,982 0.10% 0.15%
high 0.5 40,411,201,964 20,205,600,982 6.97% 10.29%
medium-high 1.0 27,515,227,962 27,515,227,962 9.49% 14.02%
medium 5.0 2,747,484,380 13,737,421,900 4.74% 7.00%
medium-low 20.0 481,935,633 9,638,712,660 3.32% 4.91%
low 50.0 97,535,696 4,876,784,800 1.68% 2.48%
bilingual very high 0.1 85,001,097,362 4,250,054,868 1.47% 0.00%
high 0.1 207,688,940,222 20,768,894,022 7.16% 0.00%
medium-high 0.2 46,777,497,823 9,355,499,565 3.23% 0.00%
medium 0.5 64,375,100,302 32,187,550,151 11.10% 0.00%
medium-low 1.0 20,361,578,347 20,361,578,347 7.02% 0.00%
low 2.0 2,503,240,829 5,006,481,658 1.73% 0.00%
very low 10.0 175,309,923 1,753,099,230 0.60% 0.00%
Appendix D Detailed Evaluation Setup
D.1 Tasks and Benchmarks
Table˜11 summarizes the tasks and benchmarks used for evaluation.
Text Classification
SIB-200 (3) and Taxi-1500 (56) are two representative multilingual topic-classification benchmarks. SIB-200 is based on Flores-200, covering 205 languages and dialects. The topics are "science/technology", "travel", "politics", "sports", "health", "entertainment", and "geography". Taxi-1500 is a sentence classification data set with 6 topics, i.e., recommendation, faith, description, sin, grace, and violence, collected from the Parallel Bible Corpus (60). It covers 1,502 typologically diverse languages spanning 112 language families.
Commonsense Reasoning
We evaluate EMMA-500 models’ commonsense-reasoning skills with two multilingual benchmarks. XCOPA (70) targets causal inference across 11 languages, XStoryCloze (52) gauges narrative completion in 11 languages.
Natural Language Inference
We evaluate on XNLI (15), where sentence pairs in different languages need to be classified as entailment, contradiction, or neutral.
Machine Translation
FLORES-200 (16) is a multilingual benchmark designed to evaluate machine translation performance across 204 language pairs involving English, yielding 408 translation directions with an emphasis on low-resource settings.
Text Summarization
We use two multilingual news summarization datasets: MassiveSumm (94) and XL-Sum (29). We subsample MassiveSumm into two sets, i.e., MassiveSumm long [^13] https://huggingface.co/datasets/MaLA-LM/MassiveSumm_long designed for longer texts, aiming for a median of 5500 tokens within a range of 3500 to 7500 tokens, and MassiveSumm short [^14] https://huggingface.co/datasets/MaLA-LM/MassiveSumm_short on shorter texts, targeting a median of 1200 tokens with a maximum of 1500 tokens. For practical reasons, we ensure a balanced number of documents across all languages, with a minimum of 100 and a maximum of 2500 documents per language in both MassiveSumm subsets. Additionally, we utilized the XL-Sum, a diverse dataset containing over one million professionally annotated article-summary pairs across 44 languages from the BBC.
Machine Reading Comprehension
We use two datasets of machine reading comprehension: BELEBELE (10) and the multilingual ARC challenge. The BELEBELE dataset (10) is a parallel multilingual machine reading comprehension benchmark spanning 122 languages across resource levels. Each of its carefully validated multiple-choice questions (with four options per question) derives from FLORES-200 passages, testing nuanced comprehension while maintaining full parallelism for cross-lingual comparison. The AI2 Reasoning Challenge (ARC) dataset (13) is a rigorously constructed benchmark for advanced question answering and reasoning evaluation. It comprises 7,787 genuine grade-school level science questions. We use the multilingual ARC translated by Lai et al. (43) using ChatGPT.
Math
We use the Multilingual Grade School Math Benchmark (MGSM) (79) that extends the GSM8K dataset (14) to evaluate cross-lingual mathematical reasoning in large language models. Derived from GSM8K’s collection of 8,500 linguistically diverse grade-school math word problems requiring multi-step reasoning, MGSM comprises 250 carefully selected problems manually translated into ten typologically diverse languages.
The evaluation benchmarks are abbreviated as follows:
  • XC: XCOPA (Cross-lingual Choice of Plausible Alternatives)
  • XSC: XStoryCloze (Cross-lingual Story Cloze Test)
  • BELE: BELEBELE, Multilingual Reading Comprehension Benchmark
  • ARC: Multilingual AI2 Reasoning Challenge
For the MGSM benchmark, we distinguish between:
  • Dir.: Direct prompting (question-only)
  • CoT: Chain-of-Thought prompting (with reasoning steps) (97)
Table 11: Evaluation statistics. Sample/Lang: average number of test samples per language; N Lang: number of languages covered.
Tasks Dataset Metric Samples/Lang N Lang Domain
Text Classification (Section˜E.1) SIB200 (3) Accuracy 204 205 Misc
Taxi1500 (56) Accuracy 111 1507 Bible
Commonsense Reasoning (Section˜E.2) XCOPA (70) Accuracy 600 11 Misc
XStoryCloze (52) Accuracy 1870 11 Misc
Natural Language Inference (Section˜E.3) XNLI (15) Accuracy 2490 15 Misc
Machine Translation (Section˜E.4) FLORES-200 (16) BLEU, chrF++ 1012 204 Misc
Summarization (Section˜E.5) XL-Sum (29) ROUGE-L, BERTScore 2537 44 News
MassiveSumm Long (94) ROUGE-L, BERTScore 3908 55 News
MassiveSumm Short (94) ROUGE-L, BERTScore 5538 88 News
Machine Comprehension (Section˜E.6) BELEBELE (10) Accuracy 900 122 Misc
ARC multilingual (43) Accuracy 1170 31 Misc
Math (Section˜E.7) MGSM direct (79) Accuracy 250 10 Misc
MGSM CoT (79) Accuracy 250 10 Misc
D.2 Evaluation Software
We evaluate benchmarks using the lm-evaluation-harness framework (26) for supported tasks, and custom or open-source scripts for others. For text classification (e.g., SIB-200, Taxi-1500), we predict the highest-probability category using next-token probabilities via Transformers (100). For translation and open-ended generation, we use vLLM (42) for faster inference.
D.3 Baselines
We compare our models with several groups of baseline approaches, covering both established and emerging LLMs.
Llama 2 series including CPT models:
  • Llama 2 base & chat (92): the second generation of Meta’s foundational 7BB parameter models release in 2023;
  • CodeLlama 2 (76): continual pre-trained model specialized for code;
  • LlaMAX Llama 2 base and Alpaca (54): continual pre-trained and instruction-tuned variants on Alpaca (87);
  • MaLA-500 Llama 2 v1 & v2 (51): multilingual adaptations with continual pre-training and low-rank adaptation (30);
  • YaYi Llama 2151515http://https://huggingface.co/wenge-research/yayi-7b-llama2: a Chinese-optimized version;
  • TowerBase and TowerInstruct Llama 2 (4): domain-specific adaptations optimized mainly for machine translation;
  • EMMA-500 Llama 2 (34): a massively multilingual variant.
Multilingual LLMs including recent advances:
Llama 3 series including CPT models:
  • Llama 3 (20): Meta’s latest generation with a cutoff or March 2023;
  • Llama 3.1 (20): Refined version with improved capabilities, which extends the data cutoff to December 2023;
  • LLaMAX Llama 3 base and Alpaca (54): continue-trained and instruction-tuned variants.
Our evaluation encompasses these representative models to ensure a comprehensive comparison. The selection covers diverse architectures and linguistic capabilities. We continue to expand our benchmarking coverage and welcome collaboration inquiries regarding additional model evaluations.
Appendix E Detailed Results
This section presents the detailed results of each evaluated task, including average scores categorized by language resource groups and per-language scores if the benchmark has a few languages. For massively multilingual benchmarks with more than a dozen of languages, we provide the results on our project website (https://mala-lm.github.io/emma-500-v2). We also provide all the model outputs of different generation tasks on the project website.
E.1 Text Classification
We evaluate with 3-shot prompting on the SIB-200 and Taxi-1500 text classification datasets. The prompt template for SIB-200 is as follows:
Topic Classification: science/technology, travel, politics, sports, health, entertainment, geography.
{examples}
The topic of the news "\${text}" is
For Taxi-1500, the prompt template is as follows:
Topic Classification: Recommendation, Faith, Description, Sin, Grace, Violence.
{examples}
The topic of the verse "\${text}" is
As for SIB-200 task, Table 12 shows that EMMA-500 Llama 3.1 8B Bi attains the highest accuracy among EMMA-500 model series (62.07%), edging its backbone by +0.65% and crossing the 60% threshold together with EMMA-500 Llama 3 8B Mono (60.62%). Yet the other two variants fall significantly below their bases: EMMA-500 Llama 3 8B Bi drops to 39.40%, while EMMA-500 Llama 3.1 8B Mono slips to 26.16%. The mixed picture—improvements for some settings, regressions for others—suggests that our current continual pre-training strategy for EMMA-500 models may not be consistently helpful for SIB-200 task, and calls for deeper analysis in terms of data mix and training strategy.
With its religious text domain, short inputs and 1500+ languages, Taxi-1500 is markedly harder. Here, the advantages of EMMA-500 are more pronounced. Where most 8 B models struggle to reach 20% accuracy, all four EMMA-500 variants hover around or above that mark. EMMA-500 Llama-3 8B-Bi achieves 25.13%, a gain of 3.40% over the base model, and EMMA-500 Llama-3.1 8B-Bi reaches 24.87%.
Table 12: 3-shot results on SIB-200 and Taxi-1500 (Accuracy %).
Model SIB-200 Taxi-1500
Avg High Med-High Medium Med-Low Low Avg High Med-High Medium Med-Low Low
Llama 2 7B 22.41 25.98 22.17 22.43 23.36 22.74 17.54 19.57 17.78 17.89 18.17 18.04
Llama 2 7B Chat 25.58 29.15 25.27 25.43 26.58 26.02 15.44 18.78 15.67 15.77 16.09 16.01
CodeLlama 7B 23.35 25.80 23.01 23.39 24.27 23.84 17.03 17.47 17.11 17.12 17.17 17.15
LLaMAX Llama 2 7B 10.61 12.83 10.96 10.66 11.11 10.78 23.52 23.08 23.54 23.55 23.52 23.53
LLaMAX Llama 2 7B Alpaca 27.89 33.36 27.69 27.88 29.74 28.66 15.09 18.87 15.31 15.43 15.70 15.57
MaLA-500 10B V1 23.25 23.00 23.44 23.03 23.30 23.13 25.27 23.53 25.08 25.00 24.83 24.76
MaLA-500 10B V2 19.30 17.83 18.57 18.42 19.51 19.04 23.39 21.27 23.35 23.07 22.94 22.72
Yayi Llama 2 7B 24.57 27.99 24.08 24.64 25.72 25.03 17.73 18.62 17.84 17.90 18.01 17.98
Tower Base Llama 2 7B V0.1 19.34 21.30 18.90 19.07 19.86 19.53 17.73 18.74 17.76 17.87 18.04 18.06
Tower Instruct Llama 2 7B V0.2 20.53 22.91 20.22 20.46 21.10 20.81 17.29 20.16 17.69 17.72 18.05 17.92
EMMA-500 Llama 2 7B 31.27 33.02 30.34 30.86 31.93 31.97 19.82 24.32 21.07 21.32 21.68 21.73
Occiglot Mistral 7B 32.69 37.77 32.91 32.87 34.07 33.27 22.26 24.74 22.53 22.61 22.69 22.62
Occiglot Mistral 7B Instruct 34.31 37.79 34.01 34.20 35.37 34.87 18.76 24.59 19.41 19.42 19.68 19.59
BLOOM 7B1 17.82 21.57 19.54 18.61 18.13 17.73 14.76 15.71 14.98 14.93 14.91 14.93
BLOOMZ 7B1 29.73 29.77 29.95 30.13 29.89 29.99 16.96 16.89 16.99 16.98 17.00 17.00
Yayi 7B 35.76 39.51 36.50 36.62 36.40 36.40 16.12 16.56 16.17 16.24 16.26 16.19
Aya 23 8B 41.50 45.45 42.17 42.43 43.42 42.20 22.64 22.82 22.63 22.56 22.56 22.52
Aya Expanse 8B 57.01 63.82 59.73 58.99 59.84 58.06 18.73 19.56 19.07 19.08 19.03 19.08
Gemma 7B 58.21 68.76 60.00 60.14 62.08 59.97 13.83 25.10 14.78 15.42 16.92 16.55
Gemma 2 9B 46.25 51.94 45.79 46.78 47.82 47.19 18.05 29.88 19.43 20.27 21.72 21.48
Qwen 1.5 7B 47.95 55.88 49.42 49.53 50.34 48.94 7.29 12.42 7.61 7.75 8.39 8.18
Qwen 2 7B 54.95 66.84 58.04 57.27 58.33 56.06 21.87 27.85 22.42 22.66 23.16 23.01
Qwen 2.5 7B 53.89 64.57 56.45 56.08 56.82 54.85 17.87 18.12 18.01 17.95 17.85 17.82
Macro-LLM GLO 7B 64.15 72.73 65.98 65.73 66.96 65.07 21.99 21.31 21.84 21.92 21.93 21.97
Llama 3 8B 63.70 73.89 66.11 66.33 67.89 65.53 21.73 32.43 22.64 23.07 23.79 23.61
Llama 3.1 8B 61.42 70.83 63.42 63.67 65.33 63.17 20.20 28.22 21.12 21.59 22.22 22.02
LLaMAX Llama 3 8B 48.60 51.41 48.19 48.94 50.16 49.70 23.01 22.32 22.84 22.85 22.82 22.91
LLaMAX Llama 3 8B Alpaca 58.97 70.50 60.57 60.96 63.71 60.91 17.71 30.56 18.56 19.32 20.83 20.48
EMMA-500 Llama 3 8B Mono 60.62 60.72 58.70 60.26 62.10 62.34 22.32 23.13 22.04 22.02 22.10 22.16
EMMA-500 Llama 3 8B Bi 39.40 37.14 39.83 38.88 38.39 39.65 25.13 29.53 26.13 26.45 26.71 26.72
EMMA-500 Llama 3.1 8B Mono 26.16 25.31 25.74 25.84 25.80 26.17 19.71 18.82 19.41 19.45 19.47 19.62
EMMA-500 Llama 3.1 8B Bi 62.07 59.80 60.15 61.60 63.31 63.33 24.87 25.71 25.50 25.45 25.40 25.42
E.2 Commonsense Reasoning
In the commonsense reasoning evaluation, all evaluations are zero-shot; accuracy is the evaluation metric. Languages are bucketed into groups according to language availability and possible corpus size, following Ji et al. (34).
Across XCOPA and XStoryCloze evaluation results (Table 13), the four EMMA-500 variants built on Llama-3/3.1 8 B clearly lift commonsense-reasoning accuracy over their backbones. The best model, EMMA-500 Llama 3.1 8B Bi, averages 67.25% on XCOPA and 68.47% on XStoryCloze—roughly +5% over vanilla Llama-3.1 and almost ahead of, all other similar size models. The performance gains are consistent across all different languages (Tables˜14 and 15). Besides, improvements concentrate in medium- and low-resource languages: EMMA-500 lifts Llama-3/3.1 by roughly +7% in medium-resource languages and flattens the usual low-resource performance drop on XStoryCloze ( 64% vs. 53% for the base), while high-resource accuracy remains competitive at around 70%.
One more interesting finding is that EMMA-500 monolingual variants are slightly behind bilingual variants by about 1%, confirming that balanced bilingual continual training adds marginal yet consistent performance gains.
Table 13: 0-shot results (Accuracy %) on commonsense reasoning: XCOPA and XStoryCloze.
Model XCOPA XStoryCloze
Avg High Medium Low Avg High Medium Low
Llama 2 7B 56.67 62.10 53.90 51.60 57.55 63.38 54.45 49.21
Llama 2 7B Chat 55.85 61.25 53.13 50.60 58.41 64.80 54.77 49.74
CodeLlama 2 7B 54.69 58.70 52.53 51.60 55.68 60.68 52.33 49.90
LLaMAX Llama 2 7B 54.38 55.50 54.13 51.40 60.36 64.34 58.82 53.47
LLaMAX Llama 2 7B Alpaca 56.60 59.80 55.17 52.40 63.83 69.08 62.01 54.33
MaLA-500 Llama 2 10B v1 53.09 53.55 53.27 50.20 53.07 58.15 49.22 48.08
MaLA-500 Llama 2 10B v2 53.09 53.55 53.27 50.20 53.07 58.15 49.22 48.08
YaYi Llama 2 7B 56.71 62.10 54.13 50.60 58.42 64.98 54.91 49.04
TowerBase Llama 2 7B 56.33 62.50 52.90 52.20 57.78 64.35 53.67 49.57
TowerInstruct Llama 2 7B 57.05 62.90 54.00 52.00 59.24 66.83 54.53 49.70
EMMA-500 Llama 2 7B 63.11 66.60 62.57 52.40 66.38 68.92 65.73 61.32
Occiglot Mistral 7B v0.1 56.67 62.80 53.37 52.00 58.10 65.18 53.28 50.03
Occiglot Mistral 7B v0.1 Instruct 56.55 62.85 52.97 52.80 59.39 66.94 54.33 50.63
BLOOM 7B 56.89 59.95 55.87 50.80 59.30 61.99 59.05 53.08
BLOOMZ 7B 54.87 56.35 54.60 50.60 57.12 61.14 55.82 49.67
YaYi 7B 56.64 59.55 55.50 51.80 60.67 64.90 59.40 52.65
Aya 23 8B 55.13 59.05 53.30 50.40 60.93 67.27 58.82 49.34
Aya Expanse 8B 56.38 61.50 53.77 51.60 64.80 73.10 61.93 49.80
Gemma 7B 63.64 70.35 61.43 50.00 65.01 69.46 64.49 54.93
Gemma 2 9B 66.33 73.40 64.27 50.40 67.67 72.47 66.69 57.64
Qwen 1.5 7B 59.44 66.85 55.90 51.00 59.85 66.62 56.04 50.56
Qwen 2 7B 60.31 68.65 56.40 50.40 61.46 69.45 56.97 50.50
Qwen 2.5 7B 61.71 72.30 56.63 49.80 62.06 69.41 58.36 51.09
Marco-LLM GLO 7B 62.45 72.20 57.77 51.60 63.87 71.26 61.23 50.66
Llama 3 8B 61.71 68.35 59.03 51.20 63.41 68.50 62.03 53.44
Llama 3.1 8B 61.71 69.30 58.80 48.80 63.58 68.66 62.09 53.87
LLaMAX Llama 3 8B 63.04 67.25 62.30 50.60 64.31 67.78 63.48 57.28
LLaMAX Llama 3 8B Alpaca 64.36 69.55 62.97 52.00 68.27 72.92 67.27 58.64
EMMA-500 Llama 3 8B Mono 66.20 69.25 66.20 54.00 67.36 69.25 66.83 63.70
EMMA-500 Llama 3 8B Bi 66.82 70.80 66.67 51.80 68.35 70.22 67.79 64.82
EMMA-500 Llama 3.1 8B Mono 65.38 68.85 65.13 53.00 67.64 69.64 67.19 63.57
EMMA-500 Llama 3.1 8B Bi 67.25 71.75 66.83 51.80 68.47 70.34 68.13 64.49
Table 14: 0-shot results (ACC %) on XCOPA in all languages
Model Avg et-acc stderr ht-acc stderr id-acc stderr it-acc stderr qu-acc stderr sw-acc stderr ta-acc stderr th-acc stderr tr-acc stderr vi-acc stderr zh-acc stderr
Llama 2 7B 56.67 48.60 2.24 50.60 2.24 62.40 2.17 65.80 2.12 51.60 2.24 52.20 2.24 53.40 2.23 56.20 2.22 54.80 2.23 62.80 2.16 65.00 2.14
Llama 2 7B Chat 55.85 47.80 2.24 50.80 2.24 62.40 2.17 67.20 2.10 50.60 2.24 52.20 2.24 50.60 2.24 55.00 2.23 55.20 2.23 61.20 2.18 61.40 2.18
CodeLlama 2 7B 54.69 46.80 2.23 51.80 2.24 57.40 2.21 63.00 2.16 51.60 2.24 48.80 2.24 55.00 2.23 55.40 2.23 53.80 2.23 55.80 2.22 62.20 2.17
LLaMAX Llama 2 7B 54.38 49.20 2.24 52.60 2.24 53.80 2.23 52.60 2.24 51.40 2.24 54.00 2.23 58.00 2.21 57.20 2.21 53.00 2.23 53.00 2.23 63.40 2.16
LLaMAX Llama 2 7B Alpaca 56.60 51.20 2.24 54.20 2.23 57.20 2.21 61.00 2.18 52.40 2.24 55.00 2.23 57.00 2.22 56.40 2.22 55.20 2.23 55.20 2.23 67.80 2.09
MaLA-500 Llama 2 10B v1 53.09 48.60 2.24 53.40 2.23 53.00 2.23 59.40 2.20 50.20 2.24 52.80 2.23 57.60 2.21 54.20 2.23 51.60 2.24 52.40 2.24 50.80 2.24
MaLA-500 Llama 2 10B v2 53.09 48.60 2.24 53.40 2.23 53.00 2.23 59.40 2.20 50.20 2.24 52.80 2.23 57.60 2.21 54.20 2.23 51.60 2.24 52.40 2.24 50.80 2.24
YaYi Llama 2 7B 56.71 48.80 2.24 50.80 2.24 62.60 2.17 67.00 2.10 50.60 2.24 53.20 2.23 55.20 2.23 54.20 2.23 55.40 2.23 63.20 2.16 62.80 2.16
TowerBase Llama 2 7B 56.33 46.00 2.23 50.20 2.24 60.20 2.19 70.80 2.04 52.20 2.24 50.60 2.24 54.40 2.23 56.00 2.22 53.80 2.23 59.20 2.20 66.20 2.12
TowerInstruct Llama 2 7B 57.05 48.80 2.24 51.60 2.24 62.00 2.17 71.00 2.03 52.00 2.24 51.00 2.24 54.20 2.23 56.40 2.22 54.60 2.23 58.60 2.20 67.40 2.10
EMMA-500 Llama 2 7B 63.11 61.40 2.18 58.00 2.21 74.20 1.96 69.40 2.06 52.40 2.24 66.20 2.12 60.00 2.19 55.60 2.22 62.00 2.17 70.20 2.05 64.80 2.14
Occiglot Mistral 7B v0.1 56.67 47.20 2.23 51.40 2.24 57.00 2.22 74.60 1.95 52.00 2.24 51.60 2.24 57.20 2.21 55.80 2.22 54.40 2.23 55.20 2.23 67.00 2.10
Occiglot Mistral 7B v0.1 Instruct 56.55 46.80 2.23 51.00 2.24 58.40 2.21 73.80 1.97 52.80 2.23 50.00 2.24 56.60 2.22 55.00 2.23 56.20 2.22 55.80 2.22 65.60 2.13
BLOOM 7B 56.89 48.20 2.24 50.80 2.24 69.80 2.06 52.80 2.23 50.80 2.24 51.80 2.24 59.20 2.20 55.40 2.23 51.00 2.24 70.80 2.04 65.20 2.13
BLOOMZ 7B 54.87 49.20 2.24 54.00 2.23 60.60 2.19 51.40 2.24 50.60 2.24 53.40 2.23 57.40 2.21 53.00 2.23 52.20 2.24 59.80 2.19 62.00 2.17
YaYi 7B 56.64 50.40 2.24 53.00 2.23 63.40 2.16 51.80 2.24 51.80 2.24 55.40 2.23 56.20 2.22 54.60 2.23 52.00 2.24 66.40 2.11 68.00 2.09
Aya 23 8B 55.13 50.20 2.24 52.20 2.24 57.00 2.22 60.40 2.19 50.40 2.24 52.20 2.24 55.60 2.22 52.60 2.24 59.60 2.20 58.80 2.20 57.40 2.21
Aya Expanse 8B 56.38 50.80 2.24 51.60 2.24 61.20 2.18 63.40 2.16 51.60 2.24 53.60 2.23 55.20 2.23 50.20 2.24 57.80 2.21 60.20 2.19 64.60 2.14
Gemma 7B 63.64 59.20 2.20 54.80 2.23 72.00 2.01 72.80 1.99 50.00 2.24 60.60 2.19 61.60 2.18 60.40 2.19 65.60 2.13 74.20 1.96 68.80 2.07
Gemma 2 9B 66.33 63.80 2.15 52.80 2.23 77.80 1.86 75.80 1.92 50.40 2.24 63.80 2.15 63.60 2.15 63.80 2.15 67.40 2.10 76.60 1.90 73.80 1.97
Qwen 1.5 7B 59.44 52.00 2.24 53.00 2.23 64.40 2.14 65.20 2.13 51.00 2.24 52.60 2.24 55.80 2.22 57.60 2.21 58.80 2.20 69.20 2.07 74.20 1.96
Qwen 2 7B 60.31 50.80 2.24 50.80 2.24 70.60 2.04 71.40 2.02 50.40 2.24 52.40 2.24 52.80 2.23 61.00 2.18 57.20 2.21 69.40 2.06 76.60 1.90
Qwen 2.5 7B 61.71 49.40 2.24 52.80 2.23 72.60 2.00 74.60 1.95 49.80 2.24 52.20 2.24 54.60 2.23 58.20 2.21 60.00 2.19 75.20 1.93 79.40 1.81
Marco-LLM GLO 7B 62.45 49.60 2.24 52.40 2.24 74.20 1.96 73.40 1.98 51.60 2.24 54.40 2.23 57.20 2.21 58.80 2.20 66.40 2.11 73.80 1.97 75.20 1.93
Llama 3 8B 61.71 53.40 2.23 52.80 2.23 71.40 2.02 71.60 2.02 51.20 2.24 58.00 2.21 60.00 2.19 58.60 2.20 62.40 2.17 71.60 2.02 67.80 2.09
Llama 3.1 8B 61.71 53.00 2.23 53.80 2.23 71.60 2.02 72.60 2.00 48.80 2.24 55.40 2.23 61.20 2.18 57.80 2.21 61.80 2.18 72.40 2.00 70.40 2.04
LLaMAX Llama 3 8B 63.04 61.80 2.18 56.40 2.22 73.00 1.99 68.40 2.08 50.60 2.24 62.60 2.17 60.80 2.19 59.20 2.20 61.40 2.18 70.40 2.04 68.80 2.07
LLaMAX Llama 3 8B Alpaca 64.36 63.00 2.16 54.00 2.23 73.00 1.99 71.80 2.01 52.00 2.24 63.40 2.16 63.20 2.16 61.20 2.18 64.00 2.15 71.80 2.01 70.60 2.04
EMMA-500 Llama 3 8B Mono 66.20 70.80 2.04 61.80 2.18 75.00 1.94 73.40 1.98 54.00 2.23 67.60 2.10 61.80 2.18 60.20 2.19 64.00 2.15 73.00 1.99 66.60 2.11
EMMA-500 Llama 3 8B Bi 66.82 71.40 2.02 60.80 2.19 78.00 1.85 73.40 1.98 51.80 2.24 68.20 2.08 63.40 2.16 58.20 2.21 66.60 2.11 73.20 1.98 70.00 2.05
EMMA-500 Llama 3.1 8B Mono 65.38 65.20 2.13 59.00 2.20 76.00 1.91 73.00 1.99 53.00 2.23 68.40 2.08 62.40 2.17 59.80 2.19 62.80 2.16 74.60 1.95 65.00 2.14
EMMA-500 Llama 3.1 8B Bi 67.25 70.40 2.04 61.00 2.18 77.00 1.88 74.60 1.95 51.80 2.24 69.60 2.06 63.20 2.16 59.80 2.19 67.20 2.10 74.60 1.95 70.60 2.04
Table 15: 0-shot results (Accuracy %) on XStoryCloze in all languages
Model Avg ar-acc stderr en-acc stderr es-acc stderr eu-acc stderr hi-acc stderr id-acc stderr my-acc stderr ru-acc stderr sw-acc stderr te-acc stderr zh-acc stderr
Llama 2 7B 57.55 49.90 1.29 77.04 1.08 67.37 1.21 50.36 1.29 53.74 1.28 59.23 1.26 48.05 1.29 63.00 1.24 50.50 1.29 54.33 1.28 59.56 1.26
Llama 2 7B Chat 58.41 50.50 1.29 78.69 1.05 67.11 1.21 50.83 1.29 54.07 1.28 59.63 1.26 48.64 1.29 65.52 1.22 52.02 1.29 53.34 1.28 62.21 1.25
CodeLlama 2 7B 55.68 50.10 1.29 71.48 1.16 63.40 1.24 50.43 1.29 49.70 1.29 55.86 1.28 49.37 1.29 59.23 1.26 50.03 1.29 53.74 1.28 59.17 1.26
LLaMAX Llama 2 7B 60.36 58.90 1.27 75.51 1.11 65.25 1.23 54.47 1.28 58.17 1.27 60.62 1.26 52.48 1.29 61.22 1.25 57.18 1.27 59.30 1.26 60.82 1.26
LLaMAX Llama 2 7B Alpaca 63.83 60.36 1.26 81.47 1.00 70.68 1.17 54.86 1.28 62.14 1.25 66.45 1.22 53.81 1.28 67.44 1.21 60.16 1.26 59.30 1.26 65.45 1.22
MaLA-500 Llama 2 10B v1 53.07 48.18 1.29 73.53 1.14 62.41 1.25 49.90 1.29 47.65 1.29 47.92 1.29 46.26 1.28 54.93 1.28 48.71 1.29 52.61 1.28 51.69 1.29
MaLA-500 Llama 2 10B v2 53.07 48.18 1.29 73.53 1.14 62.41 1.25 49.90 1.29 47.65 1.29 47.92 1.29 46.26 1.28 54.93 1.28 48.71 1.29 52.61 1.28 51.69 1.29
YaYi Llama 2 7B 58.42 49.97 1.29 79.09 1.05 68.70 1.19 50.63 1.29 54.27 1.28 61.42 1.25 47.45 1.29 64.79 1.23 50.03 1.29 53.94 1.28 62.34 1.25
TowerBase Llama 2 7B 57.78 49.17 1.29 77.23 1.08 69.82 1.18 50.76 1.29 52.88 1.28 58.31 1.27 48.38 1.29 67.04 1.21 50.36 1.29 53.14 1.28 58.50 1.27
TowerInstruct Llama 2 7B 59.24 49.31 1.29 80.87 1.01 71.61 1.16 50.69 1.29 52.95 1.28 59.56 1.26 48.71 1.29 69.36 1.19 51.49 1.29 54.14 1.28 63.00 1.24
EMMA-500 Llama 2 7B 66.38 66.25 1.22 76.44 1.09 70.02 1.18 64.73 1.23 64.92 1.23 68.63 1.19 57.91 1.27 68.50 1.20 64.73 1.23 64.66 1.23 63.40 1.24
Occiglot Mistral 7B v0.1 58.10 51.29 1.29 77.37 1.08 73.40 1.14 52.08 1.29 51.49 1.29 58.64 1.27 47.98 1.29 62.94 1.24 49.83 1.29 53.14 1.28 60.89 1.26
Occiglot Mistral 7B v0.1 Instruct 59.39 52.68 1.28 79.42 1.04 74.19 1.13 53.01 1.28 52.75 1.28 60.36 1.26 48.25 1.29 65.06 1.23 50.43 1.29 53.81 1.28 63.34 1.24
BLOOM 7B 59.30 58.57 1.27 70.55 1.17 66.18 1.22 57.25 1.27 60.42 1.26 64.53 1.23 48.91 1.29 52.75 1.28 53.94 1.28 57.31 1.27 61.88 1.25
BLOOMZ 7B 57.12 56.52 1.28 73.00 1.14 64.59 1.23 51.09 1.29 57.64 1.27 55.33 1.28 48.25 1.29 52.15 1.29 52.15 1.29 58.17 1.27 59.43 1.26
YaYi 7B 60.67 61.81 1.25 74.32 1.12 69.42 1.19 56.06 1.28 63.67 1.24 62.41 1.25 49.24 1.29 52.15 1.29 53.61 1.28 57.91 1.27 66.78 1.21
Aya 23 8B 60.93 62.61 1.25 74.59 1.12 67.31 1.21 50.96 1.29 64.26 1.23 66.64 1.21 47.72 1.29 68.70 1.19 50.36 1.29 54.00 1.28 63.14 1.24
Aya Expanse 8B 64.80 69.42 1.19 80.41 1.02 74.26 1.13 51.36 1.29 68.50 1.20 72.47 1.15 48.25 1.29 73.86 1.13 51.62 1.29 55.13 1.28 67.57 1.20
Gemma 7B 65.01 60.42 1.26 80.15 1.03 70.62 1.17 57.58 1.27 64.92 1.23 67.64 1.20 52.28 1.29 70.55 1.17 62.14 1.25 63.27 1.24 65.59 1.22
Gemma 2 9B 67.67 65.25 1.23 80.15 1.03 74.26 1.13 60.16 1.26 66.91 1.21 71.34 1.16 55.13 1.28 73.73 1.13 63.73 1.24 64.79 1.23 68.96 1.19
Qwen 1.5 7B 59.85 55.39 1.28 78.16 1.06 68.30 1.20 51.89 1.29 55.66 1.28 62.87 1.24 49.24 1.29 63.27 1.24 51.69 1.29 53.94 1.28 67.97 1.20
Qwen 2 7B 61.46 59.96 1.26 78.95 1.05 69.76 1.18 52.02 1.29 57.97 1.27 64.39 1.23 48.97 1.29 69.56 1.18 51.42 1.29 54.07 1.28 69.03 1.19
Qwen 2.5 7B 62.06 62.41 1.25 77.56 1.07 69.16 1.19 51.82 1.29 58.70 1.27 67.04 1.21 50.36 1.29 69.69 1.18 51.82 1.29 55.86 1.28 68.23 1.20
Marco-LLM GLO 7B 63.87 64.73 1.23 79.75 1.03 72.14 1.15 52.28 1.29 61.02 1.26 68.70 1.19 49.04 1.29 70.42 1.17 53.67 1.28 61.55 1.25 69.29 1.19
Llama 3 8B 63.41 58.64 1.27 78.69 1.05 70.62 1.17 55.79 1.28 62.81 1.24 65.92 1.22 51.09 1.29 68.76 1.19 56.39 1.28 63.00 1.24 65.78 1.22
Llama 3.1 8B 63.58 59.10 1.27 78.16 1.06 70.81 1.17 55.33 1.28 63.27 1.24 68.03 1.20 52.42 1.29 68.63 1.19 55.92 1.28 61.15 1.25 66.58 1.21
LLaMAX Llama 3 8B 64.31 59.96 1.26 76.97 1.08 69.09 1.19 59.23 1.26 62.87 1.24 66.78 1.21 55.33 1.28 68.30 1.20 60.69 1.26 63.60 1.24 64.59 1.23
LLaMAX Llama 3 8B Alpaca 68.27 64.73 1.23 83.19 0.96 73.00 1.14 59.96 1.26 67.50 1.21 72.40 1.15 57.31 1.27 72.80 1.15 64.00 1.24 65.19 1.23 70.88 1.17
EMMA-500 Llama 3 8B Mono 67.36 64.06 1.23 76.17 1.10 69.42 1.19 66.71 1.21 66.05 1.22 68.63 1.19 60.69 1.26 69.49 1.18 65.59 1.22 67.04 1.21 67.11 1.21
EMMA-500 Llama 3 8B Bi 68.35 66.38 1.22 75.78 1.10 71.28 1.16 66.64 1.21 66.91 1.21 69.56 1.18 63.00 1.24 70.95 1.17 67.37 1.21 67.31 1.21 66.71 1.21
EMMA-500 Llama 3.1 8B Mono 67.64 64.20 1.23 75.71 1.10 70.62 1.17 66.18 1.22 67.17 1.21 69.62 1.18 60.95 1.26 70.62 1.17 66.05 1.22 65.92 1.22 67.04 1.21
EMMA-500 Llama 3.1 8B Bi 68.47 66.18 1.22 75.12 1.11 71.34 1.16 66.45 1.22 67.77 1.20 70.68 1.17 62.54 1.25 71.41 1.16 67.50 1.21 66.58 1.21 67.64 1.20
E.3 Natural Language Inference
According to the XNLI evaluation results shown in Table 16, the bilingual EMMA-500 8B variant outperform its backbone on average (45.15% vs 44.97%) and lift low-resource accuracy by roughly +4%, closing most of the gap to the stronger Gemma 2 9 B while matching Aya Expanse 8 B. The monolingual Llama-3 variant keeps pace (44.15%), but the same Llama backbone models with monolingual continual training leads to a performance drop (39.98% in EMMA-500 Llama 3.1 8B Mono), leaving the bilingual Llama-3.1 model (44.67%) as the only Llama-3.1-based EMMA-500 choice that preserves gains in low-resource languages without a drop in high-resource languages. Overall, EMMA-500 offers modest NLI benefits concentrated in low-resource languages.
Table 16: 0-shot results on XNLI (Accuracy %).
Model Avg High Medium Low
Llama 2 7B 40.19 45.26 37.72 34.97
Llama 2 7B Chat 38.58 42.77 36.75 33.87
CodeLlama 2 7B 40.19 46.27 37.29 33.86
LLaMAX Llama 2 7B 44.27 46.53 42.64 43.03
LLaMAX Llama 2 7B Alpaca 45.09 48.47 42.80 42.89
MaLA-500 Llama 2 10B v1 38.11 42.10 35.85 34.65
MaLA-500 Llama 2 10B v2 38.11 42.10 35.85 34.65
YaYi Llama 2 7B 41.28 47.32 38.41 34.94
TowerBase Llama 2 7B 39.84 46.08 36.33 34.39
TowerInstruct Llama 2 7B 40.36 47.07 36.92 33.79
EMMA-500 Llama 2 7B 45.14 46.09 44.40 44.71
Occiglot Mistral 7B v0.1 42.35 49.90 38.39 35.19
Occiglot Mistral 7B v0.1 Instruct 40.81 47.58 37.18 34.52
BLOOM 7B 41.60 45.13 39.69 38.38
BLOOMZ 7B 37.13 40.02 35.56 34.51
YaYi 7B 39.87 43.85 38.24 35.15
Aya 23 8B 43.12 48.51 41.95 34.67
Aya Expanse 8B 45.56 50.48 44.24 38.38
Gemma 7B 42.58 46.44 41.00 38.01
Gemma 2 9B 46.74 48.50 45.11 46.49
Qwen 1.5 7B 39.47 40.95 38.80 37.83
Qwen 2 7B 42.77 47.31 41.35 36.53
Qwen 2.5 7B 43.31 47.80 41.85 37.24
Marco-LLM GLO 7B 43.99 48.74 41.44 39.57
Llama 3 8B 44.97 48.82 43.84 39.56
Llama 3.1 8B 45.62 49.61 44.04 40.83
LLaMAX Llama 3 8B 44.13 46.30 42.83 42.41
LLaMAX Llama 3 8B Alpaca 45.08 48.18 42.82 43.40
EMMA-500 Llama 3 8B Mono 44.15 47.15 41.73 42.99
EMMA-500 Llama 3 8B Bi 45.15 47.20 43.64 44.07
EMMA-500 Llama 3.1 8B Mono 39.98 42.34 39.00 37.20
EMMA-500 Llama 3.1 8B Bi 44.67 46.79 42.88 44.00
Table 17: 0-shot results (Accuracy %) on XNLI in all languages.
Model Avg ar-acc stderr bg-acc stderr de-acc stderr el-acc stderr en-acc stderr es-acc stderr fr-acc stderr hi-acc stderr ru-acc stderr sw-acc stderr th-acc stderr tr-acc stderr ur-acc stderr vi-acc stderr zh-acc stderr
Llama 2 7B 40.19 35.42 0.96 42.65 0.99 47.11 1.00 36.67 0.97 55.30 1.00 40.52 0.98 50.08 1.00 37.71 0.97 42.37 0.99 34.94 0.96 36.35 0.96 37.27 0.97 33.61 0.95 36.63 0.97 36.18 0.96
Llama 2 7B Chat 38.58 34.42 0.95 37.07 0.97 43.09 0.99 38.15 0.97 50.24 1.00 39.44 0.98 44.82 1.00 35.78 0.96 42.09 0.99 34.22 0.95 33.49 0.95 36.95 0.97 33.90 0.95 38.11 0.97 36.95 0.97
CodeLlama 2 7B 40.19 33.41 0.95 37.75 0.97 47.23 1.00 37.63 0.97 54.78 1.00 44.38 1.00 49.20 1.00 35.94 0.96 46.06 1.00 33.29 0.94 35.02 0.96 38.59 0.98 33.25 0.94 40.40 0.98 35.94 0.96
LLaMAX Llama 2 7B 44.27 33.78 0.95 46.83 1.00 48.96 1.00 42.57 0.99 54.90 1.00 47.59 1.00 47.79 1.00 45.50 1.00 45.54 1.00 43.05 0.99 41.85 0.99 43.29 0.99 44.18 1.00 43.86 0.99 34.38 0.95
LLaMAX Llama 2 7B Alpaca 45.09 34.42 0.95 46.39 1.00 49.76 1.00 43.41 0.99 58.11 0.99 48.96 1.00 51.97 1.00 45.62 1.00 46.27 1.00 43.57 0.99 40.80 0.99 43.86 0.99 44.30 1.00 43.09 0.99 35.78 0.96
MaLA-500 Llama 2 10B v1 38.11 35.94 0.96 41.20 0.99 47.51 1.00 34.46 0.95 56.18 0.99 34.10 0.95 47.59 1.00 33.65 0.95 33.94 0.95 35.22 0.96 33.69 0.95 33.82 0.95 35.02 0.96 36.02 0.96 33.25 0.94
MaLA-500 Llama 2 10B v2 38.11 35.94 0.96 41.20 0.99 47.51 1.00 34.46 0.95 56.18 0.99 34.10 0.95 47.59 1.00 33.65 0.95 33.94 0.95 35.22 0.96 33.69 0.95 33.82 0.95 35.02 0.96 36.02 0.96 33.25 0.94
YaYi Llama 2 7B 41.28 34.14 0.95 42.61 0.99 48.84 1.00 37.35 0.97 56.47 0.99 45.78 1.00 51.04 1.00 39.48 0.98 46.27 1.00 35.70 0.96 35.62 0.96 39.36 0.98 33.49 0.95 37.51 0.97 35.54 0.96
TowerBase Llama 2 7B 39.84 33.90 0.95 41.37 0.99 47.87 1.00 35.26 0.96 56.35 0.99 41.69 0.99 49.44 1.00 34.54 0.95 45.94 1.00 35.02 0.96 34.78 0.95 35.74 0.96 33.37 0.95 37.19 0.97 35.22 0.96
TowerInstruct Llama 2 7B 40.36 33.65 0.95 42.93 0.99 48.84 1.00 34.98 0.96 56.95 0.99 46.51 1.00 46.43 1.00 34.74 0.95 46.27 1.00 33.94 0.95 33.90 0.95 37.87 0.97 33.53 0.95 37.35 0.97 37.47 0.97
EMMA-500 Llama 2 7B 45.14 34.78 0.95 46.27 1.00 47.07 1.00 45.86 1.00 53.78 1.00 47.07 1.00 46.87 1.00 47.59 1.00 46.55 1.00 46.18 1.00 41.97 0.99 44.86 1.00 45.98 1.00 47.03 1.00 35.22 0.96
Occiglot Mistral 7B v0.1 42.35 33.86 0.95 41.37 0.99 51.77 1.00 37.71 0.97 55.86 1.00 51.65 1.00 51.93 1.00 35.74 0.96 47.63 1.00 34.70 0.95 37.39 0.97 43.01 0.99 33.49 0.95 38.63 0.98 40.56 0.98
Occiglot Mistral 7B v0.1 Instruct 40.81 34.38 0.95 38.84 0.98 50.84 1.00 40.00 0.98 55.66 1.00 48.63 1.00 51.69 1.00 34.30 0.95 40.04 0.98 33.17 0.94 36.35 0.96 37.99 0.97 34.06 0.95 37.59 0.97 38.63 0.98
BLOOM 7B 41.60 33.85 0.67 39.92 0.69 39.78 0.69 35.37 0.68 53.97 0.70 48.82 0.71 49.80 0.71 46.51 0.70 43.03 0.70 37.88 0.69 35.05 0.67 35.09 0.67 42.20 0.70 47.39 0.71 35.35 0.68
BLOOMZ 7B 37.13 32.69 0.94 34.02 0.95 41.69 0.99 35.82 0.96 46.87 1.00 36.02 0.96 43.05 0.99 40.40 0.98 37.47 0.97 33.61 0.95 33.09 0.94 33.73 0.95 36.83 0.97 36.67 0.97 35.02 0.96
YaYi 7B 39.87 39.80 0.98 35.78 0.96 42.61 0.99 36.47 0.96 50.52 1.00 47.71 1.00 48.19 1.00 40.04 0.98 39.12 0.98 34.06 0.95 34.34 0.95 33.17 0.94 37.07 0.97 44.18 1.00 34.94 0.96
Aya 23 8B 43.12 33.90 0.95 39.36 0.98 49.36 1.00 41.12 0.99 51.57 1.00 50.36 1.00 51.16 1.00 46.67 1.00 48.67 1.00 34.50 0.95 35.90 0.96 48.39 1.00 33.61 0.95 42.25 0.99 39.96 0.98
Aya Expanse 8B 45.56 34.10 0.95 41.89 0.99 51.04 1.00 42.17 0.99 53.86 1.00 47.35 1.00 53.65 1.00 48.31 1.00 51.00 1.00 37.03 0.97 41.04 0.99 48.84 1.00 37.07 0.97 50.12 1.00 45.98 1.00
Gemma 7B 42.58 33.49 0.95 43.49 0.99 48.63 1.00 38.11 0.97 52.05 1.00 44.14 1.00 49.76 1.00 44.34 1.00 47.39 1.00 40.64 0.98 37.95 0.97 43.25 0.99 35.42 0.96 43.29 0.99 36.67 0.97
Gemma 2 9B 46.74 34.18 0.95 49.52 1.00 51.37 1.00 43.25 0.99 53.45 1.00 51.41 1.00 52.29 1.00 47.31 1.00 49.56 1.00 45.58 1.00 49.88 1.00 50.72 1.00 44.02 1.00 45.70 1.00 32.93 0.94
Qwen 1.5 7B 39.47 34.34 0.95 40.92 0.99 42.77 0.99 36.18 0.96 49.08 1.00 37.87 0.97 43.13 0.99 38.15 0.97 38.88 0.98 35.42 0.96 44.22 1.00 38.84 0.98 33.86 0.95 44.38 1.00 33.98 0.95
Qwen 2 7B 42.77 33.69 0.95 45.46 1.00 48.19 1.00 36.83 0.97 54.26 1.00 47.23 1.00 51.41 1.00 44.98 1.00 47.39 1.00 37.31 0.97 38.27 0.97 43.53 0.99 34.02 0.95 43.57 0.99 35.38 0.96
Qwen 2.5 7B 43.31 34.38 0.95 42.85 0.99 47.31 1.00 41.45 0.99 53.65 1.00 48.92 1.00 51.57 1.00 42.49 0.99 47.71 1.00 34.22 0.95 43.33 0.99 42.57 0.99 34.18 0.95 47.39 1.00 37.63 0.97
Marco-LLM GLO 7B 43.99 33.45 0.95 45.50 1.00 50.40 1.00 37.35 0.97 53.82 1.00 50.52 1.00 51.57 1.00 42.13 0.99 46.63 1.00 34.70 0.95 44.62 1.00 44.82 1.00 39.40 0.98 45.38 1.00 39.52 0.98
Llama 3 8B 44.97 33.65 0.95 45.34 1.00 50.48 1.00 39.28 0.98 55.02 1.00 49.52 1.00 50.56 1.00 47.55 1.00 49.16 1.00 38.92 0.98 46.27 1.00 48.23 1.00 33.49 0.95 49.00 1.00 38.15 0.97
Llama 3.1 8B 45.62 33.86 0.95 45.58 1.00 51.45 1.00 38.96 0.98 55.22 1.00 50.28 1.00 51.77 1.00 49.40 1.00 49.16 1.00 39.36 0.98 48.07 1.00 49.32 1.00 35.06 0.96 47.11 1.00 39.76 0.98
LLaMAX Llama 3 8B 44.13 34.26 0.95 43.82 0.99 51.29 1.00 43.21 0.99 54.70 1.00 43.65 0.99 43.78 0.99 44.14 1.00 46.27 1.00 42.45 0.99 45.58 1.00 46.55 1.00 39.20 0.98 45.02 1.00 38.11 0.97
LLaMAX Llama 3 8B Alpaca 45.08 35.42 0.96 42.57 0.99 51.45 1.00 42.61 0.99 55.06 1.00 46.91 1.00 50.28 1.00 45.10 1.00 48.80 1.00 43.73 0.99 47.23 1.00 44.78 1.00 39.24 0.98 46.43 1.00 36.59 0.97
EMMA-500 Llama 3 8B Mono 44.15 33.01 0.94 46.06 1.00 48.59 1.00 37.91 0.97 56.39 0.99 47.43 1.00 47.95 1.00 45.78 1.00 47.79 1.00 46.14 1.00 41.00 0.99 44.50 1.00 41.81 0.99 43.09 0.99 34.74 0.95
EMMA-500 Llama 3 8B Bi 45.15 33.94 0.95 46.71 1.00 50.08 1.00 44.26 1.00 53.69 1.00 48.51 1.00 48.88 1.00 48.67 1.00 48.15 1.00 45.10 1.00 45.82 1.00 44.70 1.00 41.29 0.99 43.57 0.99 33.90 0.95
EMMA-500 Llama 3.1 8B Mono 39.98 34.66 0.95 41.33 0.99 42.61 0.99 35.62 0.96 47.99 1.00 42.33 0.99 42.41 0.99 41.16 0.99 44.58 1.00 38.47 0.98 37.19 0.97 40.56 0.98 35.94 0.96 40.68 0.98 34.10 0.95
EMMA-500 Llama 3.1 8B Bi 44.67 34.22 0.95 44.58 1.00 49.84 1.00 45.14 1.00 51.85 1.00 47.83 1.00 46.71 1.00 47.91 1.00 48.96 1.00 46.55 1.00 41.73 0.99 44.06 1.00 43.73 0.99 41.41 0.99 35.54 0.96
E.4 Machine Translation
We evaluate all models for machine translation under a 3-shot prompting setup. Specifically, we use the prompt below.
Translate the following sentence from {src_lang} to {tgt_lang}
[{src_lang}]: {src_sent}
[{tgt_lang}]:
Translation quality is assessed using BLEU (66) and chrF++ (71), implemented via sacrebleu(72). BLEU is computed using the flores200 tokenizer, which allows consistent segmentation even for languages lacking explicit word boundaries; chrF++ is calculated with a word n-gram order of 2. For transparency and reproducibility, we report full metric signatures.  [^20] BLEU: nrefs:1—case:mixed—eff:no—tok:flores200—smooth:exp —version:2.4.2; chrF++: nrefs:1—case:mixed—eff:yes—nc:6—nw:2— space:no—version:2.4.2
On translations from all other languages to English (Table 18), EMMA-500 variants based on both Llama-3 8B and Llama-3.1 8B show clear, systematic gains under monolingual and bilingual continual training. Specifically, our continual training strategy boosts the BLEU/chrF++ scores of Llama-3 8B from 23.78/43.72 to 30.38/50.86 (EMMA-500 Llama 3 8B Mono, +6.6 BLEU) and further to 34.4/54.33 (EMMA-500 Llama 3 8B Bi, +10.6 BLEU); the same schedule lifts Llama-3.1 8B from 24.19/44.1 to 27.57/48.12 (EMMA-500 Llama 3.1 8B Mono, +3.4 BLEU) and 33.6/53.5 ((EMMA-500 Llama 3 8B Bi, +9.5 BLEU).
Gains are even larger when English is the source language (Table 19): EMMA-500 Llama 3 Bi more than doubles base-model BLEU (9.93 $\rightarrow$ 24.02) and ChrF++ (24.08 $\rightarrow$ 42.15), while EMMA-500 Llama 3.1 Bi shows a comparable jump (10.11 $\rightarrow$ 23.86 BLEU). Improvements are greatest in the medium-low and low-resource languages, confirming that language-balanced bilingual continual training is especially effective for non-English translations in low-resource settings.
Table 18: 3-shot results (BLEU/chrF++) on FLORES-200, from all other languages to English (X-Eng)
Model Avg High Medium-high Medium Medium-low Low
Llama 2 7B 12.93/30.32 19.98/38.72 12.90/30.87 13.13/31.00 15.02/33.00 13.69/31.26
Llama 2 7B Chat 12.28/31.72 19.12/39.31 11.96/31.89 12.24/32.08 14.32/34.15 13.05/32.61
CodeLlama 7B 10.82/28.57 17.77/37.45 10.72/29.09 10.93/29.19 12.77/31.13 11.52/29.47
LLaMAX Llama 2 7B 1.99/13.66 3.63/21.51 2.26/15.24 2.16/14.65 2.48/16.02 2.08/14.08
LLaMAX Llama 2 7B Alpaca 22.29/42.27 34.98/56.64 24.68/45.21 24.04/44.54 26.58/47.33 23.85/44.08
MaLA-500 10B V1 2.29/13.60 5.01/16.22 2.34/13.32 2.57/13.81 2.85/14.18 2.55/13.89
MaLA-500 10B V2 2.87/15.44 5.84/18.73 2.85/15.33 3.25/15.87 3.51/16.08 3.17/15.76
Yayi Llama 2 7B 12.98/31.38 19.77/39.36 12.93/31.97 13.13/32.03 15.02/33.96 13.76/32.30
Tower Base Llama 2 7B V0.1 13.74/31.47 21.68/40.54 13.81/32.18 13.97/32.28 16.05/34.31 14.50/32.44
Tower Instruct Llama 2 7B V0.2 4.81/25.43 9.75/33.70 5.29/25.73 5.18/25.97 5.85/27.70 5.24/26.32
EMMA-500 Llama 2 7B 25.37/45.78 34.84/56.32 26.72/47.21 26.93/47.47 28.89/49.76 27.34/48.00
Occiglot Mistral 7B 13.12/31.13 19.94/38.88 12.78/30.98 13.31/31.64 15.17/33.58 14.00/32.13
Occiglot Mistral 7B Instruct 11.61/31.65 16.92/39.16 11.27/31.61 11.80/32.21 13.20/33.98 12.31/32.61
BLOOM 7B1 9.57/27.84 15.83/36.50 12.72/32.18 11.77/31.02 10.88/29.80 9.96/28.51
BLOOMZ 7B1 20.22/34.74 31.74/46.29 26.57/41.03 24.76/39.25 22.39/37.03 21.06/35.63
Yayi 7B 4.82/21.36 5.04/23.09 6.25/24.58 5.69/23.82 5.20/22.08 4.72/21.25
Aya 23 8B 13.87/32.36 19.49/39.55 15.32/34.13 14.87/33.96 16.27/35.46 14.48/33.31
Aya Expanse 8B 13.12/36.86 17.35/44.58 14.77/39.30 14.54/39.15 14.81/39.83 13.61/37.85
Gemma 7B 23.79/43.68 36.37/57.74 26.72/47.06 26.34/46.77 27.78/48.45 25.47/45.65
Gemma 2 9B 23.15/38.87 34.42/51.93 23.37/38.63 24.85/40.38 27.22/43.58 25.19/41.36
Qwen 1.5 7B 15.58/35.87 24.44/46.60 16.96/38.19 16.97/38.06 18.17/39.11 16.42/37.01
Qwen 2 7B 17.39/37.61 28.37/50.71 19.49/40.81 19.53/40.66 20.66/41.61 18.40/38.90
Qwen 2.5 7B 18.95/38.89 29.32/51.08 21.04/41.75 21.21/41.80 22.11/42.65 19.97/40.11
Macro-LLM GLO 7B 25.17/44.57 35.96/56.86 28.34/48.18 28.15/47.90 28.54/48.47 26.38/45.89
Llama 3 8B 23.78/43.72 35.67/57.20 26.55/47.03 26.51/46.96 28.08/48.72 25.49/45.72
Llama 3.1 8B 24.19/44.10 36.13/57.57 26.76/47.27 26.73/47.25 28.40/49.05 25.91/46.13
LLaMAX Llama 3 8B 0.48/4.66 0.77/6.36 0.55/5.11 0.54/5.08 0.57/5.19 0.52/4.87
LLaMAX Llama 3 8B Alpaca 25.10/45.45 37.58/58.67 27.58/48.13 27.49/48.12 29.79/50.60 26.91/47.43
EMMA-500 Llama 3 8B Mono 30.38/50.86 39.52/60.25 31.90/52.47 32.44/53.10 34.13/54.81 32.55/53.16
EMMA-500 Llama 3 8B Bi 34.40/54.33 43.48/63.45 36.45/56.47 36.98/57.00 38.45/58.44 36.72/56.69
EMMA-500 Llama 3.1 8B Mono 27.57/48.12 37.34/58.82 29.18/49.85 29.50/50.33 31.20/52.20 29.51/50.34
EMMA-500 Llama 3.1 8B Bi 33.64/53.49 42.78/62.87 35.87/55.72 36.20/56.16 37.61/57.62 35.91/55.84
Table 19: 3-shot results (BLEU/chrF++) on FLORES-200, from English to all other languages (Eng-X)
Model Avg High Medium-high Medium Medium-low Low
Llama 2 7B 4.62/15.13 10.72/26.07 4.99/16.25 4.14/15.68 6.03/18.21 4.93/16.23
Llama 2 7B Chat 4.95/16.95 10.81/26.27 5.19/17.51 4.56/17.18 6.27/19.62 5.30/18.12
CodeLlama 7B 4.27/14.94 10.03/25.25 4.61/15.56 3.92/15.40 5.58/17.75 4.57/15.98
LLaMAX Llama 2 7B 0.80/7.42 2.08/13.88 1.14/8.98 0.92/8.14 1.04/8.88 0.86/7.92
LLaMAX Llama 2 7B Alpaca 12.51/28.35 26.39/45.94 16.08/32.23 14.07/30.87 16.00/34.14 13.36/30.45
MaLA-500 10B V1 0.60/6.08 1.62/10.25 0.69/6.15 0.57/6.26 0.80/7.13 0.65/6.64
MaLA-500 10B V2 0.54/6.38 1.37/10.37 0.59/6.45 0.46/6.42 0.67/7.31 0.55/6.93
Yayi Llama 2 7B 4.41/14.87 10.77/25.96 4.89/15.86 4.07/15.38 5.90/18.04 4.81/16.04
Tower Base Llama 2 7B V0.1 4.83/16.03 11.56/25.70 5.35/16.43 4.37/16.06 6.03/18.47 5.10/17.15
Tower Instruct Llama 2 7B V0.2 3.23/15.64 7.66/24.72 3.72/16.21 3.21/16.09 4.00/17.89 3.54/16.84
EMMA-500 Llama 2 7B 15.58/33.25 28.66/46.92 18.00/34.66 16.56/34.25 18.85/37.66 16.87/35.88
Occiglot Mistral 7B 4.32/16.10 10.97/25.72 5.13/16.65 4.36/16.57 5.51/18.54 4.69/17.23
Occiglot Mistral 7B Instruct 3.99/15.80 9.80/24.99 4.69/16.44 4.01/16.39 5.09/18.34 4.31/16.92
BLOOM 7B1 2.81/11.80 6.98/20.40 4.30/14.82 3.51/13.81 3.27/13.71 2.67/12.32
BLOOMZ 7B1 7.44/16.10 22.08/32.72 12.35/22.06 10.55/20.25 9.21/18.84 7.48/16.83
Yayi 7B 4.37/13.50 12.07/27.36 6.87/18.72 5.64/16.88 5.19/15.91 4.23/14.20
Aya 23 8B 6.46/16.15 10.76/23.27 8.16/18.42 6.53/17.12 7.75/18.86 6.41/16.89
Aya Expanse 8B 6.88/23.89 10.76/32.48 8.24/26.48 7.21/25.53 7.79/26.89 6.78/25.13
Gemma 7B 9.05/23.05 18.18/37.69 10.87/26.37 10.04/25.66 11.49/27.58 9.56/24.75
Gemma 2 9B 12.09/26.48 25.23/44.22 14.93/30.26 13.67/29.59 15.29/32.05 12.71/28.50
Qwen 1.5 7B 5.87/17.77 13.12/29.98 7.14/19.71 6.01/19.13 7.13/21.03 5.93/18.87
Qwen 2 7B 5.56/17.17 12.08/28.96 6.53/19.22 5.77/18.80 6.72/20.30 5.57/18.17
Qwen 2.5 7B 5.72/17.49 12.15/28.73 6.81/19.33 5.92/18.94 6.92/20.61 5.71/18.46
Macro-LLM GLO 7B 9.27/23.34 17.74/36.96 11.10/26.36 10.30/25.76 11.19/27.21 9.51/24.79
Llama 3 8B 9.93/24.08 20.79/40.03 11.76/26.96 10.97/26.59 12.58/28.89 10.51/25.84
Llama 3.1 8B 10.11/24.69 21.29/40.67 12.00/27.60 11.10/27.20 12.76/29.55 10.66/26.46
LLaMAX Llama 3 8B 0.45/4.65 1.11/7.77 0.64/5.45 0.56/5.19 0.57/5.28 0.49/4.93
LLaMAX Llama 3 8B Alpaca 11.64/26.86 25.20/45.23 14.71/30.78 13.25/29.63 14.99/32.76 12.44/28.94
EMMA-500 Llama 3 8B Mono 20.33/38.34 33.89/51.45 22.10/39.37 21.39/39.56 24.19/42.93 22.03/41.29
EMMA-500 Llama 3 8B Bi 24.02/42.15 38.25/55.19 25.73/43.49 25.32/43.74 28.32/47.00 25.99/45.22
EMMA-500 Llama 3.1 8B Mono 19.44/37.41 32.50/50.37 21.25/38.54 20.42/38.61 23.17/41.97 21.06/40.32
EMMA-500 Llama 3.1 8B Bi 23.86/42.07 37.81/54.81 25.76/43.65 25.26/43.74 28.15/46.84 25.80/45.07
E.5 Text Summarization
We conduct zero-shot text summarization evaluations on three benchmark datasets: MassiveSumm long, MassiveSumm short, and XL-Sum. The prompt template for the three benchmarks is as follows:
Document: {text}
Based on the previous text, provide a brief single summary:
In order to try to get LLM to answer in the corresponding language of the document, we use Google Translate to translate the prompt into the corresponding language.
We evaluate the generated summaries using two widely recognized metrics: ROUGE-L (50) and BERTScore (106). ROUGE-L captures lexical overlap by measuring the longest common subsequence (LCS) between the reference and generated summaries. Precision and recall are calculated as the ratio of the LCS to the generated and reference summary lengths, respectively, with an F-score used for the final metric. BERTScore, on the other hand, measures semantic similarity by comparing contextual embeddings from pre-trained language models. Specifically, we employ the bert-base-multilingual-cased [^21] https://huggingface.co/google-bert/bert-base-multilingual-cased model, ensuring compatibility across multiple languages.
Our EMMA-500 models demonstrate consistent improvements over their Llama 3 and Llama 3.1 base models in zero-shot summarization across the MassiveSumm (long/short) and XL-Sum benchmarks (Tables 20, 21, 22). Notably, enhancements in semantic quality, measured by BERTScore, are frequently observed, especially on the MassiveSumm benchmarks. Nonetheless, Llama 3/3.1 base models do not obtain impressive summarization performance on these two subsets of MassiveSumm. On the XL-Sum benchmark, our EMMA-500 Llama 3/3.1 CPT models achieve better average performance than recent advances such as Aya 23 and Gemma 2.
Table 20: Zero-short performance of MassiveSumm long set (ROUGE-L/BERTScore). Our EMMA-500 models demonstrate consistent improvements over their Llama 3 and Llama 3.1 base models.
Model Avg High Medium-high Medium Medium-low Low
Llama 2 7B 4.74/63.89 3.88/61.92 3.88/61.92 3.40/62.75 4.78/63.95 4.78/63.95
Llama 2 7B Chat 4.73/63.52 3.88/61.57 3.88/61.57 3.42/62.34 4.76/63.62 4.76/63.62
CodeLlama 7B 5.63/64.51 4.45/60.73 4.45/60.73 4.11/63.72 5.58/64.54 5.58/64.54
LLaMAX Llama 2 7B 4.56/62.69 3.78/61.26 3.78/61.26 3.29/61.00 4.60/62.73 4.60/62.73
LLaMAX Llama 2 7B Alpaca 4.61/62.76 3.76/61.42 3.76/61.42 3.28/61.22 4.65/62.81 4.65/62.81
MaLA-500 10B V1 4.39/64.50 3.59/61.82 3.59/61.82 3.11/63.73 4.41/64.58 4.41/64.58
MaLA-500 10B V2 4.37/64.66 3.62/62.01 3.62/62.01 3.11/63.82 4.40/64.73 4.40/64.73
Yayi Llama 2 7B 4.98/64.17 4.05/62.17 4.05/62.17 3.69/63.15 5.02/64.21 5.02/64.21
Tower Base Llama 2 7B V0.1 4.81/64.51 3.88/62.39 3.88/62.39 3.43/63.49 4.85/64.59 4.85/64.59
Tower Instruct Llama 2 7B V0.2 4.82/64.61 3.84/62.52 3.84/62.52 3.40/63.49 4.85/64.67 4.85/64.67
EMMA-500 Llama 2 7B 4.79/63.80 3.81/61.99 3.81/61.99 3.43/62.66 4.82/63.88 4.82/63.88
Occiglot Mistral 7B 5.14/63.95 4.37/61.66 4.37/61.66 3.93/64.35 5.15/64.05 5.15/64.05
Occiglot Mistral 7B Instruct 5.16/63.50 4.57/61.79 4.57/61.79 4.19/63.86 5.19/63.53 5.19/63.53
BLOOM 7B1 4.88/64.36 4.00/62.34 4.00/62.34 3.76/64.65 4.92/64.45 4.92/64.45
BLOOMZ 7B1 2.91/57.20 2.12/58.71 2.12/58.71 1.89/57.54 2.94/57.17 2.94/57.17
Yayi 7B 4.95/64.24 4.26/61.87 4.26/61.87 3.81/64.25 4.99/64.32 4.99/64.32
Aya 23 8B 6.33/65.94 5.54/62.67 5.54/62.67 4.69/65.29 6.34/66.01 6.34/66.01
Aya Expanse 8B 7.44/67.66 6.29/64.39 6.29/64.39 5.65/66.65 7.48/67.76 7.48/67.76
Gemma 7B 6.18/62.14 5.25/60.19 5.25/60.19 4.80/62.14 6.23/62.19 6.23/62.19
Gemma 2 9B 5.86/59.70 4.96/57.51 4.96/57.51 4.81/59.83 5.89/59.74 5.89/59.74
Qwen 1.5 7B 6.09/59.19 5.42/58.28 5.42/58.28 5.07/60.73 6.09/59.14 6.09/59.14
Qwen 2 7B 6.65/56.31 6.47/54.91 6.47/54.91 5.43/54.74 6.64/56.16 6.64/56.16
Qwen 2.5 7B 7.44/61.62 6.88/58.85 6.88/58.85 6.19/59.98 7.50/61.59 7.50/61.59
Macro-LLM GLO 7B 6.57/57.15 5.84/53.04 5.84/53.04 5.26/55.92 6.60/57.11 6.60/57.11
Llama 3 8B 5.10/55.58 4.15/54.48 4.15/54.48 4.04/57.63 5.13/55.65 5.13/55.65
Llama 3.1 8B 5.41/56.09 4.36/55.24 4.36/55.24 4.34/57.85 5.45/56.19 5.45/56.19
LLaMAX Llama 3 8B 6.00/65.90 4.74/62.84 4.74/62.84 4.57/65.35 6.03/66.00 6.03/66.00
LLaMAX Llama 3 8B Alpaca 7.62/67.64 6.13/64.77 6.13/64.77 5.92/66.58 7.69/67.77 7.69/67.77
EMMA-500 Llama 3 8B Mono 5.38/61.06 4.61/57.84 4.61/57.84 4.42/60.75 5.41/61.08 5.41/61.08
EMMA-500 Llama 3 8B Bi 5.57/59.23 4.85/57.46 4.85/57.46 4.67/60.13 5.62/59.30 5.62/59.30
EMMA-500 Llama 3.1 8B Mono 5.44/61.89 4.70/59.04 4.70/59.04 4.49/61.99 5.48/61.98 5.48/61.98
EMMA-500 Llama 3.1 8B Bi 4.78/58.47 3.70/53.79 3.70/53.79 3.90/58.36 4.81/58.48 4.81/58.48
Table 21: Zero-short performance of MassiveSumm short set (ROUGE-L/BERTScore). Our EMMA-500 models demonstrate consistent improvements over their Llama 3 and Llama 3.1 base models.
Model Avg High Medium-high Medium Medium-low Low
Llama 2 7B 7.85/65.35 6.64/62.55 6.23/62.78 5.92/64.92 7.79/65.43 7.88/65.33
Llama 2 7B Chat 9.76/67.01 8.88/64.52 8.29/64.64 7.59/66.03 9.83/67.08 9.87/66.98
CodeLlama 7B 7.59/64.83 6.19/61.81 5.88/62.10 5.54/64.49 7.52/65.03 7.60/64.93
LLaMAX Llama 2 7B 5.22/63.06 4.47/59.76 4.10/59.92 3.46/62.60 5.16/63.15 5.26/63.10
LLaMAX Llama 2 7B Alpaca 10.71/67.92 9.52/64.83 8.93/65.09 8.12/66.93 10.73/67.99 10.81/67.90
MaLA-500 10B V1 4.97/63.51 4.34/61.56 4.00/61.65 3.53/63.32 5.04/63.76 5.13/63.74
MaLA-500 10B V2 5.02/63.75 4.34/61.61 4.00/61.69 3.55/63.46 5.09/63.97 5.18/63.95
Yayi Llama 2 7B 7.80/65.24 6.91/63.23 6.46/63.33 5.96/64.80 7.84/65.34 7.87/65.21
Tower Base Llama 2 7B V0.1 8.11/65.53 6.70/62.59 6.26/62.78 5.94/64.96 8.04/65.58 8.12/65.48
Tower Instruct Llama 2 7B V0.2 10.14/67.76 8.88/65.06 8.24/65.12 7.44/66.81 10.15/67.86 10.18/67.71
EMMA-500 Llama 2 7B 8.32/65.14 6.85/62.52 6.41/62.56 6.13/64.26 8.30/65.25 8.43/65.16
Occiglot Mistral 7B 8.16/63.65 6.39/61.67 6.05/61.74 5.79/64.17 8.15/64.00 8.26/63.96
Occiglot Mistral 7B Instruct 7.82/63.79 6.72/62.31 6.37/62.13 6.05/64.25 7.72/63.74 7.84/63.71
BLOOM 7B1 6.79/62.30 5.71/58.74 5.34/58.80 4.97/61.23 6.73/62.31 6.82/62.26
BLOOMZ 7B1 3.28/29.75 3.72/39.04 3.41/36.51 2.78/32.31 3.16/29.30 3.20/29.11
Yayi 7B 8.28/65.44 7.73/65.10 7.25/64.79 6.62/65.09 8.30/65.66 8.35/65.42
Aya 23 8B 8.43/65.85 7.49/63.54 6.99/63.73 6.14/65.59 8.47/66.04 8.47/65.85
Aya Expanse 8B 9.24/67.68 8.49/65.77 7.96/65.95 7.38/67.20 9.32/67.88 9.39/67.72
Gemma 7B 8.35/62.25 7.56/60.19 7.09/60.45 6.39/62.35 8.35/62.28 8.47/62.29
Gemma 2 9B 7.86/58.11 6.36/55.20 5.86/53.92 5.95/57.87 7.85/57.99 7.97/58.11
Qwen 1.5 7B 8.49/62.70 7.88/61.33 7.40/61.00 6.87/63.25 8.44/62.60 8.50/62.57
Qwen 2 7B 8.63/56.14 7.65/53.08 7.17/53.36 6.33/55.72 8.49/55.57 8.54/55.69
Qwen 2.5 7B 9.04/58.91 9.49/60.70 8.85/60.34 7.64/60.70 9.03/58.90 9.04/58.56
Macro-LLM GLO 7B 8.10/57.45 8.09/56.96 7.51/57.15 6.87/60.07 8.06/57.25 8.10/57.29
Llama 3 8B 6.44/50.98 4.78/48.31 4.53/48.77 4.68/52.54 6.29/50.70 6.45/50.91
Llama 3.1 8B 6.77/54.96 5.31/52.26 5.02/53.04 4.91/55.77 6.67/54.79 6.83/54.97
LLaMAX Llama 3 8B 8.77/66.46 7.05/63.80 6.60/63.77 6.27/65.97 8.65/66.58 8.76/66.47
LLaMAX Llama 3 8B Alpaca 12.44/68.95 10.27/67.22 9.54/67.08 9.42/68.35 12.44/69.09 12.48/68.94
EMMA-500 Llama 3 8B Mono 7.18/63.46 6.39/61.99 5.97/62.06 5.68/64.19 7.19/63.69 7.24/63.50
EMMA-500 Llama 3 8B Bi 7.74/63.18 6.49/60.71 6.01/60.71 5.72/63.07 7.76/63.28 7.82/63.26
EMMA-500 Llama 3.1 8B Mono 7.21/63.36 6.68/63.09 6.15/61.81 5.67/62.64 7.25/63.45 7.34/63.46
EMMA-500 Llama 3.1 8B Bi 6.67/61.64 5.61/58.37 5.21/58.16 5.01/61.53 6.67/61.59 6.75/61.64
Table 22: Zero-short performance of XL-Sum (ROUGE-L/BERTScore). Our EMMA-500 models demonstrate consistent improvements over their Llama 3 and Llama 3.1 base models.
Model Avg High Medium-high Medium Medium-low Low
Llama 2 7B 7.11/66.52 9.75/64.83 9.52/65.35 8.34/65.08 7.62/65.58 7.73/65.39
Llama 2 7B Chat 8.84/68.44 12.91/67.19 12.32/67.84 10.47/67.45 9.54/67.69 9.67/67.51
CodeLlama 7B 7.15/65.74 9.89/62.99 9.66/64.20 8.38/63.72 7.71/64.74 7.83/64.68
LLaMAX Llama 2 7B 5.29/64.59 6.86/61.74 6.66/62.29 5.99/61.96 5.71/63.24 5.79/63.17
LLaMAX Llama 2 7B Alpaca 10.11/69.24 14.44/67.88 14.31/68.35 12.42/68.26 11.12/68.38 11.19/68.19
MaLA-500 10B V1 5.45/63.96 7.98/60.89 7.40/61.16 6.38/61.14 5.88/62.48 5.94/62.55
MaLA-500 10B V2 5.44/64.28 7.80/61.21 7.28/61.62 6.34/61.51 5.86/62.71 5.91/62.79
Yayi Llama 2 7B 7.98/67.21 11.64/66.06 10.91/66.54 9.41/66.17 8.59/66.63 8.71/66.32
Tower Base Llama 2 7B V0.1 7.65/67.09 10.64/65.26 10.45/65.62 8.97/65.39 8.21/66.12 8.30/65.97
Tower Instruct Llama 2 7B V0.2 8.89/68.46 12.81/67.15 12.00/67.56 10.30/67.06 9.53/67.72 9.65/67.52
EMMA-500 Llama 2 7B 8.58/67.20 11.61/66.17 11.75/66.93 10.29/67.92 9.41/67.80 9.52/67.68
Occiglot Mistral 7B 7.33/66.20 10.38/63.81 10.17/64.44 8.86/64.39 7.93/65.00 8.02/64.88
Occiglot Mistral 7B Instruct 8.31/66.96 12.53/65.19 11.44/65.98 10.08/65.16 9.04/66.17 9.11/66.04
BLOOM 7B1 6.99/64.78 8.83/61.91 8.62/63.21 7.64/62.69 7.43/63.72 7.55/63.43
BLOOMZ 7B1 11.15/69.82 20.33/67.38 19.39/68.58 15.83/68.77 12.52/69.37 12.50/69.21
Yayi 7B 12.06/69.74 19.99/67.84 19.25/68.82 16.07/68.85 13.42/69.14 13.37/68.97
Aya 23 8B 8.68/66.79 12.91/66.50 11.90/66.45 9.92/67.34 9.25/67.11 9.43/66.91
Aya Expanse 8B 10.51/68.73 14.88/67.38 14.09/67.96 12.58/69.15 11.24/69.07 11.32/68.88
Gemma 7B 6.70/64.52 9.16/62.37 9.16/62.76 8.17/64.49 7.33/64.93 7.30/64.67
Gemma 2 9B 7.38/65.45 9.82/61.04 9.83/62.66 8.91/62.78 8.01/64.14 8.07/64.11
Qwen 1.5 7B 9.58/69.13 14.33/67.56 13.58/68.14 12.04/68.02 10.42/68.34 10.44/68.16
Qwen 2 7B 10.18/69.35 15.67/68.37 14.66/68.66 12.70/68.42 11.12/68.58 11.10/68.39
Qwen 2.5 7B 10.41/69.69 15.68/69.57 14.96/69.69 12.95/70.62 11.38/70.19 11.31/69.93
Macro-LLM GLO 7B 11.46/70.41 17.22/70.74 16.65/70.83 14.31/71.32 12.65/71.03 12.62/70.76
Llama 3 8B 8.47/67.08 10.96/64.16 11.24/65.16 10.12/65.04 9.29/65.82 9.31/65.79
Llama 3.1 8B 8.57/66.97 10.97/63.60 11.19/64.55 10.14/64.81 9.45/65.66 9.43/65.66
LLaMAX Llama 3 8B 8.28/66.72 10.85/64.80 10.97/65.33 9.71/66.92 9.04/67.25 9.07/66.99
LLaMAX Llama 3 8B Alpaca 11.39/69.99 15.45/69.66 15.82/70.20 13.36/70.75 12.44/70.53 12.42/70.34
EMMA-500 Llama 3 8B Mono 9.11/66.12 12.96/64.51 13.25/65.23 11.30/66.81 10.09/66.83 9.96/66.50
EMMA-500 Llama 3 8B Bi 9.10/66.72 13.30/65.28 13.11/65.91 11.17/67.34 10.10/67.46 9.94/67.12
EMMA-500 Llama 3.1 8B Mono 9.66/67.21 12.88/65.35 13.18/66.30 11.53/67.78 10.63/67.93 10.58/67.66
EMMA-500 Llama 3.1 8B Bi 8.56/65.90 12.19/64.10 11.75/64.40 10.51/66.43 9.51/66.53 9.47/66.31
E.6 Machine Reading Comprehension
We first evaluate on the BELEBELE dataset (10), a comprehensive multilingual benchmark for machine reading comprehension, spanning 122 languages across both high- and low-resource categories. Notably, this benchmark demonstrates significant difficulty, with even the English subset posing substantial challenges to state-of-the-art language models like Llama 3. As shown in Table˜23, which presents zero-shot performance across language resource groups, our continual pre-training approach on Llama 3 and 3.1, as well as LLaMAX CPT models, yields some degraded improvements over the base Llama model. However, when measured by the number of languages with improved performance, our EMMA-500 CPT is better than the base models. Contemporary models such as Aya Expanse, Llama 3.1, and Qwen 2 demonstrate more competent performance, suggesting advances in multilingual understanding and reasoning capacities. On the harder ARC-Multilingual set, the bilingual Llama-3 variant essentially ties the base model ($\approx$ 35%), while the others dip by up to one point. These results suggest that EMMA-style bilingual continual pre-training alone is insufficient for cross-language passage‐question reasoning at the 8 B scale, the fundamental limitations of Llama models, and the average-based evaluation practice.
Table 23: 0-shot results (Accuracy %) on BELEBELE, and 5-shot results (Accuracy %) on the multilingual ARC.
Model BELEBELE ARC Multilingual
Avg High Med-high Medium Med-low Low Avg High Medium Low
Llama 2 7B 26.27 26.76 26.35 26.07 26.41 26.27 27.56 33.12 27.31 21.02
Llama 2 7B Chat 29.05 31.84 29.97 28.95 29.47 29.09 28.02 33.69 27.79 21.29
CodeLlama 2 7B 27.38 27.37 27.33 27.30 27.30 27.38 25.23 28.86 24.64 21.65
LLaMAX Llama 2 7B 23.09 23.23 23.15 23.07 23.10 23.08 26.09 30.00 25.92 21.48
LLaMAX Llama 2 7B Alpaca 24.48 25.46 24.82 24.41 24.60 24.49 31.06 36.89 31.85 22.49
MaLA-500 Llama 2 10B v1 22.96 23.02 22.98 22.97 22.98 22.97 21.16 21.92 20.48 21.32
MaLA-500 Llama 2 10B v2 22.96 23.02 22.98 22.97 22.98 22.97 21.16 21.92 20.48 21.32
YaYi Llama 2 7B 28.32 29.64 28.67 28.11 28.37 28.26 28.40 34.30 28.35 21.11
TowerBase Llama 2 7B 26.36 27.43 26.85 26.29 26.48 26.34 27.94 35.32 26.82 20.51
TowerInstruct Llama 2 7B 27.93 29.88 28.51 27.57 28.19 27.92 30.10 38.88 28.85 21.16
EMMA-500 Llama 2 7B 26.75 28.32 28.18 27.58 27.14 26.94 29.53 34.10 29.82 23.34
Occiglot Mistral 7B v0.1 30.16 32.25 30.94 30.02 30.40 30.15 29.77 38.39 28.51 21.03
Occiglot Mistral 7B v0.1 Instruct 32.05 34.14 32.62 31.74 32.40 32.08 30.88 40.29 29.65 21.13
BLOOM 7B 24.11 24.25 24.52 24.12 24.11 24.08 23.65 26.27 22.72 21.89
BLOOMZ 7B 39.32 45.43 43.67 41.51 40.08 39.51 23.95 26.94 22.74 22.18
YaYi 7B 37.97 44.37 42.71 40.49 38.72 38.09 24.44 27.96 23.29 21.91
Aya 23 8B 40.08 43.85 41.71 39.22 40.93 39.81 31.08 40.05 30.02 21.61
Aya Expanse 8B 46.98 52.22 48.99 46.57 48.36 46.93 36.56 47.87 36.15 23.09
Gemma 7B 43.37 52.63 47.83 44.82 45.43 43.94 38.68 46.46 40.47 26.06
Gemma 2 9B 54.49 64.10 58.90 55.83 56.85 55.05 44.15 54.59 46.18 27.82
Qwen 1.5 7B 41.83 48.86 44.79 42.18 43.00 41.78 28.93 35.55 28.14 21.92
Qwen 2 7B 49.31 57.62 52.20 50.04 51.16 49.48 33.82 43.88 32.64 23.17
Qwen 2.5 7B 54.11 63.47 57.91 55.47 56.24 54.30 35.30 46.43 34.31 23.00
Marco-LLM GLO 7B 53.95 63.54 58.67 55.50 56.20 54.31 36.34 46.72 35.88 24.11
Llama 3 8B 40.73 46.07 42.92 41.03 41.87 40.88 34.80 42.43 35.53 24.06
Llama 3.1 8B 45.19 52.50 48.01 45.65 46.74 45.34 34.93 42.43 35.89 24.00
LLaMAX Llama 3 8B 36.96 40.06 37.92 36.78 37.68 37.09 33.54 39.81 34.84 23.59
LLaMAX Llama 3 8B Alpaca 39.41 44.06 41.10 39.52 40.53 39.61 34.53 41.34 35.56 24.34
EMMA-500 Llama 3 8B Mono 39.73 43.65 41.40 39.96 40.76 40.01 33.22 38.32 34.56 24.67
EMMA-500 Llama 3 8B Bi 39.84 43.51 41.09 40.05 40.74 39.99 34.84 40.37 36.45 25.31
EMMA-500 Llama 3.1 8B Mono 38.86 41.68 39.34 38.49 39.62 38.94 34.00 39.38 35.39 25.01
EMMA-500 Llama 3.1 8B Bi 37.00 40.59 38.36 37.21 37.94 37.25 34.59 39.85 36.21 25.38
E.7 Math
We use the Multilingual Grade School Math Benchmark (MGSM) (79) to evaluate the mathematical reasoning in LLMs. As shown in Table˜24, we evaluate model performance using 3-shot prompting with flexible answer matching. Our assessment employs two distinct prompting strategies: (1) Direct: direct question prompting, and (2) Chain-of-Thought (CoT): prompting accompanied by step-by-step reasoning examples (97). The results show that our CPT models, as well as LLaMAX CPT models, obtain decreased average performance on direct prompting. Our EMMA-500 Llama 3.1 8B Mono model slightly improves the base model. Notably, our EMMA-500 Llama 3/3.1 CPT models significantly improve the performance on low-resource languages for both direct and CoT prompting.
Tables˜25 and 26 present the per-language performance across all the languages by direct and CoT prompting.
Table 24: 3-shot results (Accuracy %) on MGSM obtained by direct and CoT prompting. Our EMMA-500 CPT models significantly improve the performance on low-resource languages.
Model Direct CoT
Avg High Medium Low Avg High Medium Low
Llama 2 7B 6.69 8.07 2.13 1.20 6.36 7.60 2.13 0.80
Llama 2 7B Chat 10.22 13.73 2.13 0.80 10.91 13.53 2.80 1.60
CodeLlama 2 7B 5.93 7.07 2.93 1.20 6.64 8.73 2.67 2.00
LLaMAX Llama 2 7B 3.35 4.00 2.00 0.80 3.62 4.33 2.27 2.40
LLaMAX Llama 2 7B Alpaca 5.05 5.20 4.00 1.60 6.35 8.07 4.13 0.80
MaLA-500 Llama 2 10B v1 0.91 1.33 0.27 0.00 0.73 1.27 0.00 0.00
MaLA-500 Llama 2 10B v2 0.91 1.33 0.27 0.00 0.73 1.27 0.00 0.00
YaYi Llama 2 7B 7.09 9.47 1.47 1.20 7.22 8.73 2.27 1.20
TowerBase Llama 2 7B 6.15 8.33 1.73 0.80 6.16 8.60 2.40 0.80
TowerInstruct Llama 2 7B 7.24 9.53 1.73 2.00 8.24 10.47 1.87 1.20
EMMA-500 Llama 2 7B 17.02 19.20 11.87 2.40 18.09 20.00 13.20 2.80
Occiglot Mistral 7B v0.1 13.31 16.87 4.53 1.60 14.07 18.80 3.60 1.20
Occiglot Mistral 7B v0.1 Instruct 22.76 29.80 7.47 2.80 22.16 30.40 7.87 2.80
BLOOM 7B 2.87 2.60 2.80 3.60 2.29 2.20 1.47 2.00
BLOOMZ 7B 2.55 2.67 2.40 1.20 2.15 1.67 3.07 2.00
YaYi 7B 2.76 2.93 1.47 2.40 3.02 2.93 2.40 2.00
Aya 23 8B 22.29 30.67 3.47 0.80 24.71 35.07 5.47 2.40
Aya Expanse 8B 43.02 55.47 18.93 5.20 41.45 54.67 18.53 5.20
Gemma 7B 38.22 36.60 38.27 27.20 35.78 34.67 37.07 26.80
Gemma 2 9B 32.95 28.00 35.73 30.80 44.69 36.07 52.00 47.20
Qwen 1.5 7B 31.56 40.00 16.00 4.00 30.36 40.60 14.80 2.40
Qwen 2 7B 48.95 54.40 38.80 14.80 51.47 58.93 39.07 14.40
Qwen 2.5 7B 53.78 65.33 36.13 8.00 55.60 68.87 36.40 8.80
Marco-LLM GLO 7B 51.85 63.80 32.80 11.60 52.02 62.93 34.93 11.60
Llama 3 8B 27.45 27.87 26.13 5.60 28.13 28.53 26.67 5.20
Llama 3.1 8B 28.36 29.00 26.13 4.40 27.31 27.27 25.47 8.40
LLaMAX Llama 3 8B 20.80 22.73 20.40 6.80 19.96 20.93 22.40 3.60
LLaMAX Llama 3 8B Alpaca 14.18 14.87 11.87 6.00 17.16 20.87 12.27 4.00
EMMA-500 Llama 3 8B Mono 23.53 22.33 25.60 12.00 25.33 24.67 27.73 13.60
EMMA-500 Llama 3 8B Bi 23.49 23.00 24.67 10.80 26.29 24.87 30.67 11.60
EMMA-500 Llama 3.1 8B Mono 24.95 23.67 27.20 11.20 27.35 26.60 30.27 10.80
EMMA-500 Llama 3.1 8B Bi 23.85 23.67 25.33 11.20 25.76 26.27 27.87 12.80
Table 25: 3-shot results (Accuracy %) on MGSM in all languages by direct prompting and flexible matching.
Model Avg bn bn-stderr de de-stderr en en-stderr es es-stderr fr fr-stderr ja ja-stderr ru ru-stderr sw sw-stderr te te-stderr th th-stderr zh zh-stderr
Llama 2 7B 6.69 2.80 1.05 8.00 1.72 17.60 2.41 11.20 2.00 12.00 2.06 2.40 0.97 8.00 1.72 2.80 1.05 1.20 0.69 0.80 0.56 6.80 1.60
Llama 2 7B Chat 10.22 2.80 1.05 16.80 2.37 22.80 2.66 19.60 2.52 19.20 2.50 2.40 0.97 14.40 2.22 0.80 0.56 0.80 0.56 2.80 1.05 10.00 1.90
CodeLlama 2 7B 5.93 1.60 0.80 8.80 1.80 12.80 2.12 8.80 1.80 10.40 1.93 5.60 1.46 6.00 1.51 2.00 0.89 1.20 0.69 5.20 1.41 2.80 1.05
LLaMAX Llama 2 7B 3.35 2.80 1.05 3.60 1.18 6.00 1.51 2.00 0.89 7.20 1.64 3.20 1.12 2.40 0.97 1.60 0.80 0.80 0.56 1.60 0.80 5.60 1.46
LLaMAX Llama 2 7B Alpaca 5.05 4.00 1.24 3.60 1.18 10.80 1.97 6.00 1.51 6.40 1.55 3.20 1.12 4.40 1.30 4.80 1.35 1.60 0.80 3.20 1.12 7.60 1.68
MaLA-500 Llama 2 10B v1 0.91 0.00 0.00 0.00 0.00 1.20 0.69 2.00 0.89 2.40 0.97 1.20 0.69 2.00 0.89 0.40 0.40 0.00 0.00 0.40 0.40 0.40 0.40
MaLA-500 Llama 2 10B v2 0.91 0.00 0.00 0.00 0.00 1.20 0.69 2.00 0.89 2.40 0.97 1.20 0.69 2.00 0.89 0.40 0.40 0.00 0.00 0.40 0.40 0.40 0.40
YaYi Llama 2 7B 7.09 3.20 1.12 8.40 1.76 15.60 2.30 16.00 2.32 10.40 1.93 5.20 1.41 5.60 1.46 0.80 0.56 1.20 0.69 0.40 0.40 11.20 2.00
TowerBase Llama 2 7B 6.15 2.40 0.97 8.40 1.76 11.60 2.03 9.20 1.83 8.80 1.80 4.80 1.35 10.00 1.90 1.20 0.69 0.80 0.56 1.60 0.80 8.80 1.80
TowerInstruct Llama 2 7B 7.24 1.60 0.80 10.00 1.90 15.20 2.28 15.60 2.30 12.80 2.12 1.60 0.80 10.00 1.90 1.60 0.80 2.00 0.89 2.00 0.89 7.20 1.64
EMMA-500 Llama 2 7B 17.02 8.80 1.80 23.20 2.68 34.00 3.00 28.00 2.85 25.60 2.77 9.20 1.83 22.80 2.66 16.80 2.37 2.40 0.97 10.00 1.90 6.40 1.55
Occiglot Mistral 7B v0.1 13.31 3.20 1.12 21.20 2.59 30.00 2.90 27.20 2.82 21.60 2.61 6.40 1.55 15.20 2.28 2.40 0.97 1.60 0.80 8.00 1.72 9.60 1.87
Occiglot Mistral 7B v0.1 Instruct 22.76 4.80 1.35 34.00 3.00 46.40 3.16 40.00 3.10 31.60 2.95 18.40 2.46 23.60 2.69 6.40 1.55 2.80 1.05 11.20 2.00 31.20 2.94
BLOOM 7B 2.87 2.40 0.97 1.60 0.80 4.00 1.24 3.60 1.18 1.20 0.69 2.00 0.89 3.60 1.18 4.00 1.24 3.60 1.18 2.00 0.89 3.60 1.18
BLOOMZ 7B 2.55 3.20 1.12 1.60 0.80 3.60 1.18 2.40 0.97 3.20 1.12 2.80 1.05 3.60 1.18 2.80 1.05 1.20 0.69 1.20 0.69 2.40 0.97
YaYi 7B 2.76 2.40 0.97 2.00 0.89 6.00 1.51 2.40 0.97 5.60 1.46 1.60 0.80 1.20 0.69 1.20 0.69 2.40 0.97 0.80 0.56 4.80 1.35
Aya 23 8B 22.29 3.20 1.12 37.20 3.06 50.00 3.17 41.20 3.12 39.20 3.09 27.60 2.83 34.00 3.00 2.80 1.05 0.80 0.56 4.40 1.30 4.80 1.35
Aya Expanse 8B 43.02 22.40 2.64 69.20 2.93 78.40 2.61 72.00 2.85 63.60 3.05 50.80 3.17 69.60 2.92 12.80 2.12 5.20 1.41 21.60 2.61 7.60 1.68
Gemma 7B 38.22 34.40 3.01 44.80 3.15 58.80 3.12 48.00 3.17 39.60 3.10 16.80 2.37 41.20 3.12 37.60 3.07 27.20 2.82 42.80 3.14 29.20 2.88
Gemma 2 9B 32.95 29.60 2.89 40.40 3.11 56.40 3.14 50.80 3.17 36.00 3.04 1.20 0.69 35.20 3.03 37.60 3.07 30.80 2.93 40.00 3.10 4.40 1.30
Qwen 1.5 7B 31.56 12.40 2.09 44.00 3.15 55.20 3.15 48.40 3.17 45.20 3.15 16.80 2.37 43.20 3.14 7.20 1.64 4.00 1.24 28.40 2.86 42.40 3.13
Qwen 2 7B 48.95 44.40 3.15 65.60 3.01 80.80 2.50 76.00 2.71 69.60 2.92 1.60 0.80 67.20 2.98 16.00 2.32 14.80 2.25 56.00 3.15 46.40 3.16
Qwen 2.5 7B 53.78 26.80 2.81 63.20 3.06 83.20 2.37 74.00 2.78 66.40 2.99 52.40 3.16 72.80 2.82 20.00 2.53 8.00 1.72 61.60 3.08 63.20 3.06
Marco-LLM GLO 7B 51.85 31.60 2.95 64.40 3.03 77.60 2.64 71.20 2.87 63.20 3.06 50.40 3.17 63.60 3.05 18.00 2.43 11.60 2.03 48.80 3.17 70.00 2.90
Llama 3 8B 27.45 17.60 2.41 39.60 3.10 50.80 3.17 47.20 3.16 36.40 3.05 3.60 1.18 37.60 3.07 24.00 2.71 5.60 1.46 36.80 3.06 2.80 1.05
Llama 3.1 8B 28.36 20.00 2.53 41.20 3.12 55.20 3.15 48.40 3.17 38.40 3.08 3.60 1.18 40.00 3.10 20.80 2.57 4.40 1.30 37.60 3.07 2.40 0.97
LLaMAX Llama 3 8B 20.80 13.20 2.15 20.40 2.55 24.40 2.72 26.00 2.78 22.00 2.63 20.80 2.57 22.80 2.66 21.20 2.59 6.80 1.60 26.80 2.81 24.40 2.72
LLaMAX Llama 3 8B Alpaca 14.18 12.00 2.06 17.60 2.41 25.20 2.75 17.20 2.39 17.20 2.39 11.20 2.00 15.60 2.30 10.80 1.97 6.00 1.51 12.80 2.12 10.40 1.93
EMMA-500 Llama 3 8B Mono 23.53 20.00 2.53 30.00 2.90 36.00 3.04 34.80 3.02 30.00 2.90 3.20 1.12 33.20 2.98 33.20 2.98 12.00 2.06 23.60 2.69 2.80 1.05
EMMA-500 Llama 3 8B Bi 23.49 21.20 2.59 30.80 2.93 35.60 3.03 36.80 3.06 34.00 3.00 2.80 1.05 30.40 2.92 30.00 2.90 10.80 1.97 22.80 2.66 3.20 1.12
EMMA-500 Llama 3.1 8B Mono 24.95 21.20 2.59 34.80 3.02 39.60 3.10 35.60 3.03 33.20 2.98 1.20 0.69 34.80 3.02 36.80 3.06 11.20 2.00 23.60 2.69 2.40 0.97
EMMA-500 Llama 3.1 8B Bi 23.85 20.80 2.57 38.00 3.08 33.20 2.98 39.60 3.10 32.40 2.97 2.00 0.89 28.40 2.86 27.20 2.82 11.20 2.00 28.00 2.85 1.60 0.80
Table 26: 3-shot results (Accuracy %) on MGSM in all languages by CoT prompting and flexible matching.
Model Avg bn bn-stderr de de-stderr en en-stderr es es-stderr fr fr-stderr ja ja-stderr ru ru-stderr sw sw-stderr te te-stderr th th-stderr zh zh-stderr
Llama 2 7B 6.36 2.00 0.89 7.60 1.68 16.00 2.32 12.40 2.09 9.20 1.83 3.60 1.18 8.00 1.72 2.00 0.89 0.80 0.56 1.60 0.80 8.00 1.72
Llama 2 7B Chat 10.91 2.40 0.97 17.60 2.41 27.20 2.82 19.20 2.50 18.80 2.48 4.00 1.24 15.20 2.28 2.00 0.89 0.80 0.56 2.80 1.05 11.60 2.03
CodeLlama 2 7B 6.64 1.60 0.80 9.20 1.83 13.20 2.15 10.80 1.97 10.80 1.97 4.40 1.30 7.60 1.68 1.20 0.69 1.60 0.80 6.00 1.51 4.00 1.24
LLaMAX Llama 2 7B 3.62 3.60 1.18 3.60 1.18 7.20 1.64 3.60 1.18 5.20 1.41 2.80 1.05 2.80 1.05 0.80 0.56 0.80 0.56 2.00 0.89 4.80 1.35
LLaMAX Llama 2 7B Alpaca 6.35 3.20 1.12 4.80 1.35 15.20 2.28 9.60 1.87 5.20 1.41 4.40 1.30 3.60 1.18 4.80 1.35 0.40 0.40 4.80 1.35 6.80 1.60
MaLA-500 Llama 2 10B v1 0.73 0.00 0.00 0.40 0.40 0.40 0.40 0.40 0.40 1.60 0.80 1.60 0.80 2.40 0.97 0.40 0.40 0.00 0.00 0.00 0.00 0.80 0.56
MaLA-500 Llama 2 10B v2 0.73 0.00 0.00 0.40 0.40 0.40 0.40 0.40 0.40 1.60 0.80 1.60 0.80 2.40 0.97 0.40 0.40 0.00 0.00 0.00 0.00 0.80 0.56
YaYi Llama 2 7B 7.22 2.80 1.05 8.40 1.76 16.80 2.37 12.40 2.09 10.40 1.93 4.80 1.35 7.20 1.64 2.00 0.89 1.20 0.69 3.20 1.12 12.40 2.09
TowerBase Llama 2 7B 6.16 3.60 1.18 8.00 1.72 11.20 2.00 8.40 1.76 7.60 1.68 3.60 1.18 8.40 1.76 1.20 0.69 0.80 0.56 2.80 1.05 9.20 1.83
TowerInstruct Llama 2 7B 8.24 1.20 0.69 12.80 2.12 18.80 2.48 15.20 2.28 12.00 2.06 2.40 0.97 13.60 2.17 1.60 0.80 1.20 0.69 3.20 1.12 10.80 1.97
EMMA-500 Llama 2 7B 18.09 8.40 1.76 22.40 2.64 37.60 3.07 25.60 2.77 21.60 2.61 8.00 1.72 22.80 2.66 21.20 2.59 2.40 0.97 11.60 2.03 16.40 2.35
Occiglot Mistral 7B v0.1 14.07 2.40 0.97 21.60 2.61 33.20 2.98 26.80 2.81 22.40 2.64 6.40 1.55 17.20 2.39 3.20 1.12 1.20 0.69 6.00 1.51 11.20 2.00
Occiglot Mistral 7B v0.1 Instruct 22.16 4.00 1.24 33.20 2.98 43.20 3.14 42.00 3.13 31.60 2.95 16.80 2.37 24.40 2.72 5.60 1.46 1.60 0.80 8.40 1.76 24.80 2.74
BLOOM 7B 2.29 2.00 0.89 1.60 0.80 4.80 1.35 2.40 0.97 2.40 0.97 0.40 0.40 4.40 1.30 1.60 0.80 2.00 0.89 2.00 0.89 2.40 0.97
BLOOMZ 7B 2.15 2.00 0.89 2.00 0.89 2.40 0.97 3.20 1.12 3.20 1.12 2.00 0.89 1.60 0.80 2.40 0.97 2.00 0.89 1.60 0.80 1.20 0.69
YaYi 7B 3.02 4.80 1.35 3.20 1.12 4.80 1.35 3.60 1.18 4.00 1.24 2.00 0.89 1.20 0.69 1.60 0.80 2.80 1.05 0.80 0.56 6.00 1.51
Aya 23 8B 24.71 6.00 1.51 40.80 3.11 48.80 3.17 41.20 3.12 40.00 3.10 28.40 2.86 37.20 3.06 5.60 1.46 2.40 0.97 8.40 1.76 6.80 1.60
Aya Expanse 8B 41.45 21.20 2.59 68.80 2.94 78.80 2.59 76.80 2.68 65.20 3.02 29.60 2.89 67.60 2.97 11.20 2.00 2.40 0.97 20.00 2.53 2.80 1.05
Gemma 7B 35.78 36.40 3.05 40.80 3.11 60.80 3.09 45.20 3.15 42.00 3.13 3.60 1.18 39.20 3.09 38.80 3.09 24.80 2.74 44.00 3.15 4.80 1.35
Gemma 2 9B 44.69 48.80 3.17 53.60 3.16 72.00 2.85 60.40 3.10 50.00 3.17 2.00 0.89 50.00 3.17 53.60 3.16 44.40 3.15 55.20 3.15 1.60 0.80
Qwen 1.5 7B 30.36 12.80 2.12 43.20 3.14 58.40 3.12 50.40 3.17 42.00 3.13 11.20 2.00 36.80 3.06 8.80 1.80 4.00 1.24 25.60 2.77 26.00 2.78
Qwen 2 7B 51.47 45.20 3.15 66.00 3.00 82.00 2.43 75.20 2.74 71.60 2.86 2.80 1.05 68.40 2.95 17.20 2.39 17.60 2.41 58.40 3.12 60.80 3.09
Qwen 2.5 7B 55.60 25.60 2.77 69.60 2.92 81.60 2.46 73.60 2.79 68.00 2.96 60.00 3.10 74.80 2.75 17.20 2.39 9.20 1.83 56.80 3.14 72.80 2.82
Marco-LLM GLO 7B 52.02 32.00 2.96 65.20 3.02 78.80 2.59 70.00 2.90 64.00 3.04 48.80 3.17 64.40 3.03 18.00 2.43 9.60 1.87 56.00 3.15 63.60 3.05
Llama 3 8B 28.13 19.20 2.50 38.40 3.08 54.00 3.16 46.00 3.16 36.00 3.04 2.00 0.89 38.00 3.08 25.20 2.75 6.40 1.55 39.20 3.09 4.00 1.24
Llama 3.1 8B 27.31 19.20 2.50 42.00 3.13 52.80 3.16 42.80 3.14 34.80 3.02 2.80 1.05 37.20 3.06 25.20 2.75 5.60 1.46 35.60 3.03 1.60 0.80
LLaMAX Llama 3 8B 19.96 12.40 2.09 20.40 2.55 24.80 2.74 25.60 2.77 18.40 2.46 18.80 2.48 20.40 2.55 23.20 2.68 3.60 1.18 30.80 2.93 19.60 2.52
LLaMAX Llama 3 8B Alpaca 17.16 14.80 2.25 18.00 2.43 26.00 2.78 22.40 2.64 24.00 2.71 13.60 2.17 22.40 2.64 12.40 2.09 7.20 1.64 11.60 2.03 13.20 2.15
EMMA-500 Llama 3 8B Mono 25.33 23.20 2.68 34.00 3.00 37.60 3.07 40.40 3.11 32.80 2.98 1.60 0.80 33.60 2.99 32.40 2.97 12.80 2.12 24.00 2.71 2.80 1.05
EMMA-500 Llama 3 8B Bi 26.29 22.00 2.63 34.40 3.01 40.40 3.11 42.80 3.14 37.20 3.06 0.80 0.56 31.60 2.95 33.60 2.99 12.00 2.06 29.20 2.88 1.20 0.69
EMMA-500 Llama 3.1 8B Mono 27.35 22.00 2.63 38.00 3.08 40.00 3.10 41.60 3.12 36.40 3.05 3.20 1.12 37.20 3.06 42.40 3.13 10.40 1.93 26.40 2.79 2.00 0.89
EMMA-500 Llama 3.1 8B Bi 25.76 24.40 2.72 38.00 3.08 35.20 3.03 38.80 3.09 35.20 3.03 1.60 0.80 33.20 2.98 35.20 3.03 8.80 1.80 25.20 2.75 1.20 0.69
Appendix F Ethics Considerations
Data Collection and Use
This work involves the compilation and use of large-scale monolingual and bilingual textual corpora covering over 500 languages. All datasets used for continual pre-training are obtained from publicly available sources or established linguistic resources that permit academic research use. Given the scale and diversity of the data, we acknowledge the potential presence of sensitive or harmful content. Future work could include filtering mechanisms or annotation efforts to improve the quality and safety of multilingual data.
Representation of Low-Resource Languages
Our research aims to expand LLM coverage to underrepresented languages. However, the quantity and quality of data for low-resource languages remain uneven. This may inadvertently lead to biased model behaviors or inadequate performance for some linguistic groups.
Community Involvement
This work does not directly involve community collaboration. However, we recognize the importance of inclusive research practices and welcome future partnerships with linguists, native speakers, and regional institutions to improve the quality and cultural relevance of multilingual language technologies.